### v1

# Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting

Abstract
Time series forecasting task predicts future trends based on historical information. Recent U-Net-based methods have demonstrated superior performance in predicting real-world datasets. However, the performance of these models is lower than patch-based models or linear models. In this work, we propose a symmetric and hierarchical framework, Kernel-U-Net, which cuts the input sequence into slices at each layer of the network and then computes them using kernels. Furthermore, it generalizes the concept of convolutional kernels in classic U-Net to accept custom kernels that follow the same design pattern. Compared to the existing linear or transformer-based solution, our model contains 3 advantages: 1) A small number of parameters: the parameters size is $O(log(L)^{2})$ where $L$ is the look-back window size, 2) Flexibility: its kernels can be customized and fitted to the datasets, 3) Computation efficiency: the computation complexity of transformer modules is reduced to $O(log(L)^{2})$ if they are placed close to the latent vector. Kernel-U-Net accuracy was greater than or equal to the state-of-the-art model on six (out of seven) real-world datasets.
1 Introduction
Time series forecasting predicts future trends based on recent historical information. It allows experts to track the incoming situation and react timely in critical cases. Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city Lai et al. (2018), monitoring influenza-like illness weekly patient cases Wu et al. (2021), monitoring electricity transformer temperature in the electric power long-term deployment Zhou et al. (2021) or forecasting temperature, pressure and humidity in weather station Liu et al. (2022b) etc.
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods (e.g., ARIMA Khandelwal et al. (2015) and machine learning techniques (e.g., GBRT Persson et al. (2017)) to deep learning-based solutions, such as RNN recurrent neural networks Tokgöz and Ünal (2018), LSTM Kong et al. (2017), Temporal Convolutional Network (TCN) Hewage et al. (2020) and Transformer-based model Li et al. (2019).
Large-scale time series forecasting addresses the problem of predicting a longer future horizon with a larger look-back window. With the increasing size of data, the cost of learning deep models increases crucially. Moreover, models such as LSTM or RNN suffer gradient vanishing problems in learning long sequences. To balance the quality of prediction and efficiency of learning deep models, researchers proposed hierarchical structure in Duan et al. (2022), pyramidal structure in Pyraformer Liu et al. (2022a), AutoEncoders Asres et al. (2022), U-Net-like model in Yformer Madhusudhanan et al. (2023) or patches in PatchTST Nie et al. (2023). Meanwhile, there is an increasing trend in combining different components for learning particular time series. For example in time series classification, researchers studied hybrid models such as combing GRU Elsayed et al. (2019), LSTM Lin et al. (2020); Khan et al. (2021) or Transformer Huang et al. (2020b) with linear or CNN layers Hewage et al. (2020).
Among the Transformer models applying to time series data, Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022) are the best variants that improve the quality of prediction. Meanwhile, a recent paper Zeng et al. (2023) gives a simple linear model NLinear which outperforms pure transformer-based models on a variety of common benchmarks and challenges the Transformer models for time series forecasting. Later, the authors in PatchTST Nie et al. (2023) argued that applying transformer modules on a point-wise sequence degrades the performance. By adding a linear patch layer, they successfully relieved the overfitting problem of transformer modules and they reached state-of-the-art result on several large datasets.
We remark that models are particularly sensible to different types of datasets. For example, NLinear is more robust against overfitting in small-size datasets such as Etth and EttmZeng et al. (2023) and PatchTST is more expressive in large-size datasets such as Weather and Traffic etc.Nie et al. (2023). At the same time, it is also important to notice the parameter size in Linear models such as NLinear is in $O(L^{2})$ and the computation complexity is $O(L)$. Meanwhile, the parameter size in a transformer block is $O(1)$ but the computational complexity is $O(L^{2})$ in general Transformer modules. It is reduced to $O(Llog(L))$ using Propsparse and to $O(\frac{L}{S}log(\frac{L}{S}))$ using patches, where $S$ is the size of the patch. Based on previous analysis, there is a motivation to propose a more flexible framework for combining the advantages of different modules in high-paced development and testing.
As a classical and highly expressive model in medical image segmentation, Convolutional U-net has an elegant symmetrical encoder and decoder structureRonneberger et al. (2015). It coincides well with the time series forecasting problem because the inputs and outputs are drawn from the same distribution. The first U-Net-like model for time series forecasting was YformerMadhusudhanan et al. (2023), which used Transformers in the encoder and decoder. As mentioned before, applying a Transformer on point-wise data can lead to overfitting. Therefore, we are interested in answering the question of whether there is a U-net-like structure that performs well for the time series forecasting task.
Figure 1: Structure of Kernel U-Net
We propose a general hierarchical and symmetrical framework, Kernel-U-Net, inspired by convolutional U-net Ronneberger et al. (2015) and Yformer Madhusudhanan et al. (2023) for time series forecasting. Kernel-U-Net offers $O(log(L)^{2})$ parameters size for processing a time series of size $L$. Meanwhile, it provides convenience for composing particular models with custom kernels that follow the same design pattern. Since the hierarchical structure reduces the length of the sequence at each level exponentially, the complexity of learning such sequences is simultaneously reduced. In particular, the computation complexity of Transformer and Long Short-term Memory(LSTM) modules is reduced to $O(log(L)^{2})$ and $O(log(L))$ if they are only employed in the layers close to the latent vector. Therefore, Kernel-U-Net helps reduce the computation cost of the Transformer and relieves the gradient vanishing problem in LSTM in learning long sequences.
To fully study the performance and efficiency of Kernel-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. Our results show that in time series forecasting, Kernel-U-Net meets or outperforms baseline models, such as NLinear and PatchTST, in the majority cases.
In summary, the contributions of this work include:
  • We propose Kernel-U-Net, a general hierarchical and symmetrical framework that partitions input sequences into small slices and learns them at each layer of the network.
  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the same design pattern. The parameter size is guaranteed in $O(log(L)^{2})$.
  • The computation complexity of Transformer module is $O(Llog(L))$ and can be reduced to $O(log(L)^{2})$ if it is only employed in the layers close to the latent vector.
  • Kernel-U-Net achieves results with MSE close to or better than the state-of-the-art methods on 6 (out of 7) time series data sets.
We conclude that Kernel-U-Net is a highly potential candidate for large-scale time series forecasting tasks. It is a general hierarchical and symmetrical framework with low complexity in terms of parameters. It reaches state-of-the-art results in the majority cases. With flexible customized kernels, it adapts well in high-paced development and test context.
2 Related works
Transformer
Transformer Vaswani et al. (2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks. It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route. Bidirectional Encoder Representations from Transformers (BERT)Devlin et al. (2019) is designed to pre-train deep bidirectional representations from the unlabeled text by joint conditioning on both left and right context in all layers. Generative Pre-trained Transformer 3(GPT-3) pre-trained a massive transformer-based model using 175 billion parameters Brown et al. (2020) and achieved strong performance on different types of downstream natural language tasks without requiring any fine-tuning.
Later, Vision Transformers (ViTs) Dosovitskiy et al. (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al. (2009). Swin Transformer Liu et al. (2021) proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans Li et al. (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. It proposes also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer Zhou et al. (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. It proposes a ProbSparse self-attention mechanism to efficiently replace the canonical self-attention. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. PyraformerLiu et al. (2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass.
Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing. Autoformer Wu et al. (2021) introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods. Frequency Enhanced Decomposed Transformer (FEDformer)Zhou et al. (2022) captures global properties of time series with seasonal-trend decomposition. The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
Recently, YFormer Madhusudhanan et al. (2023) proposed a U-Net-like architecture that combines coarse and fine-grained features to capture long-range effects on all scale levels. PatchTST Nie et al. (2023) adopted the patches method from ViTs and proposed a time series transformer with a linear patch layer. It reaches state-of-the-art results on many time series forecasting benchmarks. The author claims Reduction in time and space complexity of learning transformer block from $O(L^{2})$ to $O((\frac{L}{S})^{2})$, where $L$ is sequence size and $S$ is stride size.
U-Net
U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al. (2015). U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is reshaped into small slices and reduced by kernels step by step into a latent vector. At the decoder stage, the vector is developed by kernels for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al. (2015), biomedical 3D-Image segmentation Çiçek et al. (2016), time series segmentation Perslev et al. (2019), image super-resolution Han et al. (2022) and image denoising Zhang et al. (2023). Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al. (2016), 1D-U-Net Azar et al. (2022), Weighted Res-UNet Xiao et al. (2018) with residual structure and weighted attention mechanism, U-Net++ Zhou et al. (2018) with nested U-net inside each layer, UNet3+ Huang et al. (2020a) that allows full-scale skip connections incorporating low-level details with high-level semantics from feature maps in different scales and Swin-U-Net Cao et al. (2023) that only use transformer blocks at each layer of U-Net. Modification on U-Net layers has been studied by Selective Kernel U-Net Byra et al. (2020) which merges information from different sizes of convolutional kernels and Large Kernel U-Net Jia et al. (2022) which combines TransMorph layers for better performance in medical image segmentation.
In time series processing, U-time Perslev et al. (2019) is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer Madhusudhanan et al. (2023) is the first U-Net based model for time series forecasting task. In particular, YFormer applied transformer blocks on each layer of U-Net and capitalized on multi-resolution feature maps effectively.
Hierarchical and hybrid model
In time series processing, learning sequences of length $L$ requires $O(L)$ computations for recurrent models such as RNN, LSTM, GRU, and $O(L^{2})$ computations for transformer-based models. With ProbSparse Attention, the complexity can be reduced to $O(Llog(L))$. To balance the expressivity of complex modules and computational efficiency, researchers investigated hierarchical architecture and hybrid methods that merge different modules into the network.
For example, authors in Du et al. (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al. (2017) applied hierarchical LSTM and GRU was used for document classification. In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification, authors in LogTrans Li et al. (2019), Informer Zhou et al. (2021)) and Pyraformer Liu et al. (2022a) studied hierarchical or pyramidal transformers for time series forecasting.
Kernel-U-Net is a general hierarchical and symmetrical framework with low complexity in terms of parameter size. By replacing the convolutional kernel in classical U-Net, our model allows users to define custom kernels such as Attention, LSTM, or simple linear multiplication with a non-linear activation function for extracting features on complex time series. There are 3 advantages in Kernel-U-Net: 1) A small number of parameters: the total parameter of model is in order of $O(log(L)^{2})$, 2) Flexibility: the kernels can be replaced easily by arbitrary function that follows the same design mode so that non-linear kernels can be easily combined into it, 3) Computation efficiency: the computation complexity of transformer modules is reduce to $O(log(L)^{2})$ if they placed close to latent vector.
Algorithm 1 Kernel Wrapper
Input: kernel $f_{i}$, input segments $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
Output: Instance of Kernel Wrapper
1:  # Initiation of main kernel function $f$ is
2:  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
3:  # Holds variables for global operation
4:  _unet_skip_output = None
5:  _unet_skip_input = None
6:
7:  # Define the forward function as :
8:  Reshape $x$ into (-1, $L_{in}$,$M_{in}$)
9:  if _unet_skip_input is not None then
10:     $x=x+\_unet\_skip\_input$
11:  end if
12:  # Apply $f$ on $x$
13:  $y=f_{i}(x)$
14:  _unet_skip_output = $y$
15:  return $y$
3 Method
3.1 Problem Formulation
Let us note by $x\in R^{M\times N}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the size of the feature. Let L be the length of memory or the look-back window, so $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$ is a slice of length $L$ of all features. It contains historical information about the system at instant $t$. We call $X\in R^{N\times M\times L}$ the trajectory matrix which represents the sliced historical data for each instant $t\in[0,N-L-1]$.
Time series forecasting.
In the case of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$. Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window. Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps. Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
Algorithm 2 Kernel U-Net Encoder
Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net Encoder
1:  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
2:  # Create first layer
[…]
4:  # Create intermediate layers
[…]
6:     layers.append(KernelWrapper($f_{i},L_{i},M_{i},1,M_{i+1}$))
7:  end for
8:  # Create last layer
9:  layers.append(KernelWrapper($f_{n+m},L_{n+m},M_{n+m},$ $L_{h},M_{h}$))
10:  encoder = nn.Sequential (*layers)
3.2 Kernel U-Net
Kernel U-Net follows the hierarchical structure of convolutional U-Net. It accepts a custom kernel to replace the linear kernel inside the original convolutional U-Net. More precisely, it repeatedly applies custom kernels on small slices of large-scale input data. In this context, Kernel U-Net avoids problems in processing long sequences such as gradient vanishing in recurrent models and quadradic complexity in transformers-based models. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence, and the creation of kernel wrapper, encoder, decoder, and U-net.
Hierarchical Partition of the input sequence.
Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step. Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small slices $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all slices is equal to its original trajectory matrix:
$X_{t}=Concat([[X_{t,i,j}\forall i\in[1,\frac{L}{L_{1}}]]\forall j\in[1,\frac{M}{M_{1}}]])$.
The partitioned sequences are processed by kernels in the encoder. Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated segments $\hat{X}_{t+L,i,j}$.
Hierarchical processing with kernels in the encoder and decoder
Let us consider $X$ as a batch of trajectory matrix $X\in R^{B\times L\times M}$, where $B$ is the batch size. The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage. By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder.
At the stage of the encoder, the shape of input $X$ is $(B,L,M)$. Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem. We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of slices and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$. After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer. Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector.
At the stage of the decoder, the operations are reversed. We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels. At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$.
The algorithm is optimized in implementation because the for loop is expensive in the forward function of a PyTorch module. In our implementation, we nest the reshape and transpose operations in the kernel wrapper to avoid iteration over kernels. The complete algorithm is described in the next section.
Algorithm 3 Kernel U-Net Decoder
Input: $L_{h}$, $M_{h}$, $[L_{n},...,L_{2}]$, $[L_{n+m},...,L_{n+2}]$, $L_{1}$, $M_{1}$, $[f_{n+m},...,f_{n+2},$ $f_{n},...,f_{2}]$, $[M_{n+m},...,M_{n+2},$ $M_{n},...,M_{2}]$, $L_{1}$, $M_{1}$.
Output: Instance of Kernel U-Net Decoder
1:  # Create a list of multiples
[…]
3:  # Create first layer
4:  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$ $L_{n+m},$ $M_{n+m}$)]
5:  # Create intermediate layers
6:  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
[…]
8:  end for
9:  # Create last layer
10:  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
11:  decoder = nn.Sequential (*layers)
Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns. A kernel wrapper is an interface that executes the given kernel inside the encoder and decoder. It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector. By default, a kernel reduces the length of a segment from $L_{i}$ to $1$ in the encoder and develops the length of a segment from $1$ to $L_{i}$ in the decoder. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
Algorithm 4 Kernel U-Net
Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net
1:  # Initiation of encoder and decoder in K-U-N.
2:  encoder = KernelUNetEncoder(inputs)
3:  decoder = KernelUNetDecoder(inputs)
4:
5:  # Define the forward function as
6:  z = encoder(x)
7:  for  $i$ in $[1,...,n]$ do
8:     output = encoder.layers[i]._unet_skip_output
9:     decoder.layers[$n-i+1$]. _unet_skip_input = output
10:  end for
11:  y= decoder(z)
12:  return y
Creation of Kernel U-Net
Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of segment-unit $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$.
Remark that the hidden dimension of the model is the size of intermediate output vectors within layers. It corresponds to channel size in a convolutional network. This dimension can be augmented for a larger passage of information. The kernel wrapper is described in algorithm 1. The creation of Kernel U-Net Encoder, Decoder, and complete U-Net are described in algorithms 2, 3, 4.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.369 0.404 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713
192 0.397 0.426 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792
336 0.429 0.428 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809
720 0.433 0.45 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865
ETTh2
96 0.271 0.336 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525
192 0.332 0.376 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931
336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835
720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625
ETTm1
96 0.285 0.337 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571
192 0.327 0.367 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669
336 0.357 0.392 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871
720 0.408 0.422 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823
ETTm2
96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453
192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563
336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887
720 0.348 0.378 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338
Electricity
96 0.129 0.225 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368
192 0.148 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386
336 0.168 0.263 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394
720 0.205 0.299 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439
Traffic
96 0.360 0.252 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391
192 0.377 0.263 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379
336 0.389 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42
720 0.427 0.290 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472
Weather
96 0.146 0.200 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384
192 0.192 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544
336 0.241 0.287 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523
720 0.310 0.340 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741
Table 1: Multivariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
3.3 Custom Kernels
Linear kernel
The linear kernel is a simple matrix multiplication. Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
$Z_{i}=f_{i}(X)=X*w+b$
, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector. Remark that linear Kernel U-Net is an equivalent version to convolutional U-Net whose stride size is $L_{i}$. We also test a non-linear kernel that takes a non-linear activation function such as Tanh in the intermediate hidden layers.
Transformer kernel
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.Wen et al. (2023)
Positional encoding (PE).
In a vanilla Transformer, the input sequence is added by a positional encoding vector given by:
$PE(t)=\left\{\begin{array}[]{ll}sin(\omega it)&ifi\%2=0\\ cos(\omega it)&ifi\%2=1\\ \end{array}\right.$
, where $\omega i$ is the frequency of waves generated at each dimension. By adding positional information to the input, the transformer can consider the same patterns that occur in different time steps as different patterns.
Multiple head attentions.
With the Query-Key-Value $(QKV)$ model, the scaled dot-product attention used by Transformer is given by
$Attention(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{D_{k}}}))V$
, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values. Remark that in the context of self-attention, $Q=K=V=inputdata$.
The multiple attention block contains repeated heads and can be given by
$MultiHeadAttn(Q,K,V)=Concat(h_{1},...,h_{H})W^{O}$
, where $h_{i}=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}).$
LSTM kernel
The Long Short-term memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$. Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$. The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$. The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state. The formulation of LSTM is as follows:
$f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$
$i_{t}=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$
$C^{\prime}_{t}=tanh(W_{C}[h_{t-1},x_{t}]+b_{C})$
$C_{t}=f_{t}C_{t-1}+i_{t}C^{\prime}_{t}$
$o_{t}=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$
$h_{t}=o_{t}tanh(C_{t})$
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer    LogTrans
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.052 0.176 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468
192 0.067 0.202 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409
336 0.074 0.22 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546
720 0.074 0.223 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629
ETTh2
96 0.114 0.262 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379
192 0.149 0.304 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429
336 0.158 0.319 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437
720 0.178 0.34 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387
ETTm1
96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171
192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317
336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459
720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579
ETTm2
96 0.061 0.181 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208
192 0.085 0.22 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275
336 0.104 0.25 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302
720 0.149 0.303 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321
Table 2: Univariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
The LSTM Kernel used in our experiment contains an LSTM cell and a linear layer that takes all hidden states of a sequence as input and gives a vector as output. The experimental result is shown in the section 4
4 Experiments and Results
Datasets.
We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Traffic and Electricity. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
Here, we followed the experiment setting in Zeng et al. (2023) and partitioned the data into 12, 4, and 4 months for training, validation, and testing respectively for the ETT dataset. The data is split into [0.7,0.1,0.2] for training, validation, and testing for dataset Weather, Traffic, and Electricity.
Baselines and Experimental Settings.
In the time series forecasting task, we follow the experiment setting in NLinear Zeng et al. (2023), which take $L={336,720}$ step historical data as input and forecast $T\in{96,192,336,720}$ step value in the future. We replace the last value normalization with mean value normalization for univariate forecasting in the ETT dataset. We include five recent methods: PatchTSTNie et al. (2023), NLinear, DLinearZeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021)and LogTrans Li et al. (2019). We merge the result reported in Nie et al. (2023) and Zeng et al. (2023) for taking their best in a supervised setting.
Following previous works Wu et al. (2021), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the core metrics to compare performance for Forecasting problems. We take a look back at Windows 336 and 720 in experiments. The list of multiples is respectively [3, 4, 4, 5] and [4, 4, 7] and the segment-unit input length is 3. The input dimension is 1 as we follow the channel-independent setting in Zeng et al. (2023). The hidden dimensions are 128 or 256 for multivariate tasks and 5 or 30 for univariate tasks depending on of dataset. We use the Random Erase Zhong et al. (2020) from image processing for data augmentation. The learning rate is selected in [0.00005, 0.001] depending on the dataset. The training epoch is 300 and the patience is 30 in general.
Instance Normalization and Mean Normalization Instance normalization Ulyanov et al. (2017) Kim et al. (2021) contains a step of substruction with the average value of the input sequence and a division of its standard deviation. It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model. Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises in the small dataset. We followed the mean normalization as the preprocessing step and compared it with instance normalization for selected models.
Model Variants.
We propose 4 kernels and 5 versions of kernel U-Net for experiments. The kernels tested are Linear, Linear with 1 hidden layer and non-linear activation, Transformer, and LSTM.
  • Linear U-Net is a baseline method that contains a list of linear kernels.
  • Linear-1-Hidden U-Net replaces 2 kernels close to latent vector to kernels with hidden layer and nonlinear activation.
  • Linear-5-Hidden U-Net replaces all linear kernels with non-linear kernels.
  • Transformer U-Net replaces 2 kernels close to latent vector to transformer kernels and keeps the rest as linear kernels.
  • LSTM U-Net replaces 2 kernels close to latent vector to LSTM kernels and keeps the rest as linear kernels.
The result shows that the performance of Kernel U-Net reaches or outperforms the state-of-art methods in univariate setting (Table 2) and reaches similar or better results in multivariate setting (Table 1).
5 Conclusion
In this paper, we propose Kernel U-Net, a symmetric and hierarchical framework that can effectively process long temporal dependencies with $O(log(L)^{2})$ number of parameters. Specifically, Kernel U-Net cut input sequences into small slices at each layer in the network for applying customized kernels.
In the future, we hope to develop more kernels that adapt well to particular datasets. We also hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.

### v2

@@
-Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting
+Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting
 Abstract
 Time series forecasting task predicts future trends based on historical information.
-Recent U-Net-based methods have demonstrated superior performance in predicting real-world datasets.
-However, the performance of these models is lower than patch-based models or linear models.
-In this work, we propose a symmetric and hierarchical framework, Kernel-U-Net, which cuts the input sequence into slices at each layer of the network and then computes them using kernels.
-Furthermore, it generalizes the concept of convolutional kernels in classic U-Net to accept custom kernels that follow the same design pattern.
-Compared to the existing linear or transformer-based solution, our model contains 3 advantages: 1) A small number of parameters: the parameters size is $O(log(L)^{2})$ where $L$ is the look-back window size, 2) Flexibility: its kernels can be customized and fitted to the datasets, 3) Computation efficiency: the computation complexity of transformer modules is reduced to $O(log(L)^{2})$ if they are placed close to the latent vector.
-Kernel-U-Net accuracy was greater than or equal to the state-of-the-art model on six (out of seven) real-world datasets.
+Transformer-based U-Net architectures, despite their success in medical image segmentation, have limitations in both expressiveness and computation efficiency in time series forecasting as evidenced in YFormer.
+To tackle these challenges, we introduce Kernel-U-Net, a symmetric and hierarchical U-shape neural network architecture.
+The kernel-U-Net encoder compresses gradually input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series.
+Specifically, Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels.
+Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear.
+Experiments on seven real-world datasets, considering both multivariate and univariate settings, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model PatchTST in the majority of cases and outperforms Yformer.
+The source code for Kernel-U-Net will be made publicly available for further research and application.
 1 Introduction
@@
 (2022b) etc.
-Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods (e.g., ARIMA Khandelwal et al.
-(2015) and machine learning techniques (e.g., GBRT Persson et al.
-(2017)) to deep learning-based solutions, such as RNN recurrent neural networks Tokgöz and Ünal (2018), LSTM Kong et al.
+Over the past few decades, time series forecasting solutions have evolved from traditional statistical methodsKhandelwal et al.
+(2015) and machine learning techniquesPersson et al.
+(2017) to deep learning-based solutions, such as recurrent neural networks(RNN) Tokgöz and Ünal (2018), Long Short-term Memory (LSTM) Kong et al.
 (2017), Temporal Convolutional Network (TCN) Hewage et al.
 (2020) and Transformer-based model Li et al.
-(2019).
-Large-scale time series forecasting addresses the problem of predicting a longer future horizon with a larger look-back window.
-With the increasing size of data, the cost of learning deep models increases crucially.
-Moreover, models such as LSTM or RNN suffer gradient vanishing problems in learning long sequences.
-To balance the quality of prediction and efficiency of learning deep models, researchers proposed hierarchical structure in Duan et al.
-(2022), pyramidal structure in Pyraformer Liu et al.
-(2022a), AutoEncoders Asres et al.
-(2022), U-Net-like model in Yformer Madhusudhanan et al.
-(2023) or patches in PatchTST Nie et al.
-(2023).
-Meanwhile, there is an increasing trend in combining different components for learning particular time series.
-For example in time series classification, researchers studied hybrid models such as combing GRU Elsayed et al.
-(2019), LSTM Lin et al.
-(2020); Khan et al.
-(2021) or Transformer Huang et al.
-(2020b) with linear or CNN layers Hewage et al.
 (2019).
@@
 (2021), and FEDformer Zhou et al.
-(2022) are the best variants that improve the quality of prediction.
-Meanwhile, a recent paper Zeng et al.
-(2023) gives a simple linear model NLinear which outperforms pure transformer-based models on a variety of common benchmarks and challenges the Transformer models for time series forecasting.
-Later, the authors in PatchTST Nie et al.
-(2023) argued that applying transformer modules on a point-wise sequence degrades the performance.
-By adding a linear patch layer, they successfully relieved the overfitting problem of transformer modules and they reached state-of-the-art result on several large datasets.
-We remark that models are particularly sensible to different types of datasets.
-For example, NLinear is more robust against overfitting in small-size datasets such as Etth and EttmZeng et al.
-(2023) and PatchTST is more expressive in large-size datasets such as Weather and Traffic etc.Nie et al.
-(2023).
-At the same time, it is also important to notice the parameter size in Linear models such as NLinear is in $O(L^{2})$ and the computation complexity is $O(L)$.
-Meanwhile, the parameter size in a transformer block is $O(1)$ but the computational complexity is $O(L^{2})$ in general Transformer modules.
-It is reduced to $O(Llog(L))$ using Propsparse and to $O(\frac{L}{S}log(\frac{L}{S}))$ using patches, where $S$ is the size of the patch.
-Based on previous analysis, there is a motivation to propose a more flexible framework for combining the advantages of different modules in high-paced development and testing.
-As a classical and highly expressive model in medical image segmentation, Convolutional U-net has an elegant symmetrical encoder and decoder structureRonneberger et al.
-(2015).
-It coincides well with the time series forecasting problem because the inputs and outputs are drawn from the same distribution.
-The first U-Net-like model for time series forecasting was YformerMadhusudhanan et al.
-(2023), which used Transformers in the encoder and decoder.
-As mentioned before, applying a Transformer on point-wise data can lead to overfitting.
-Therefore, we are interested in answering the question of whether there is a U-net-like structure that performs well for the time series forecasting task.
-Figure 1: Structure of Kernel U-Net
-We propose a general hierarchical and symmetrical framework, Kernel-U-Net, inspired by convolutional U-net Ronneberger et al.
-(2015) and Yformer Madhusudhanan et al.
-(2023) for time series forecasting.
-Kernel-U-Net offers $O(log(L)^{2})$ parameters size for processing a time series of size $L$.
-Meanwhile, it provides convenience for composing particular models with custom kernels that follow the same design pattern.
-Since the hierarchical structure reduces the length of the sequence at each level exponentially, the complexity of learning such sequences is simultaneously reduced.
-In particular, the computation complexity of Transformer and Long Short-term Memory(LSTM) modules is reduced to $O(log(L)^{2})$ and $O(log(L))$ if they are only employed in the layers close to the latent vector.
-Therefore, Kernel-U-Net helps reduce the computation cost of the Transformer and relieves the gradient vanishing problem in LSTM in learning long sequences.
-To fully study the performance and efficiency of Kernel-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets.
-Our results show that in time series forecasting, Kernel-U-Net meets or outperforms baseline models, such as NLinear and PatchTST, in the majority cases.
+(2022) are the best variants that incrementally improved the quality of prediction.
+As a recent paper Zeng et al.
+(2023) challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in Nie et al.
+(2023) argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches.
+By adding a linear patch layer, their model PatchTST successfully relieved the overfitting problem of transformer modules and reached state-of-the-art results.
+We observe that models display distinct strengths depending on the dataset type.
+For instance, NLinear stands out for its efficiency in handling univariate time series tasks, particularly with small-size datasets.
+On the other hand, PatchTST is noteworthy for its expressiveness in multivariate time series tasks on large-size datasets.
+These contrasting attributes highlight the necessity for a unified but flexible architecture.
+This architecture would aim to integrate various modules easily, allowing for specific customized solutions.
+Such integration should not only ensure a balance between computational efficiency and expressiveness but also respond to requests for rapid development and testing.
+The Convolutional U-net, as a classic and highly expressive model in medical image segmentationRonneberger et al.
+(2015), features a symmetric encoder and decoder structure that is elegant in its design.
+This model’s structure is particularly suited to the time series forecasting task, as both inputs and outputs in this context are typically derived from the same distribution.
+The first U-shape model adapted for time series forecasting was the YformerMadhusudhanan et al.
+(2023), which incorporated Transformers in both its encoder and decoder components.
+As mentioned previously, employing Transformers on point-wise data has the potential to cause overfitting issues.
+Therefore, our investigation aims to discover if there is a U-shape architecture effective in time series forecasting, which also possesses the capability to integrate various modules flexibly, thus facilitating the specific customization of solutions.
+To tackle this problem, we propose a symmetric and hierarchical architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net and Yformer for time series forecasting.
+K-U-Net provides convenience for composing particular models with custom kernels that follow the design pattern.
+By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data.
+Furthermore, the hierarchical structure of K-U-Net exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning such sequences.
+Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
+To fully study the performance and efficiency of K-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets.
+We compose 30 variants of K-U-Net by placing different kernels at different layers and then we choose the best candidates for fine-tuning.
+Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinearZeng et al.
+(2023) and PatchTST, in the majority of cases.
 In summary, the contributions of this work include:
-  • We propose Kernel-U-Net, a general hierarchical and symmetrical framework that partitions input sequences into small slices and learns them at each layer of the network.
-  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the same design pattern.
-The parameter size is guaranteed in $O(log(L)^{2})$.
-  • The computation complexity of Transformer module is $O(Llog(L))$ and can be reduced to $O(log(L)^{2})$ if it is only employed in the layers close to the latent vector.
-  • Kernel-U-Net achieves results with MSE close to or better than the state-of-the-art methods on 6 (out of 7) time series data sets.
-We conclude that Kernel-U-Net is a highly potential candidate for large-scale time series forecasting tasks.
-It is a general hierarchical and symmetrical framework with low complexity in terms of parameters.
-It reaches state-of-the-art results in the majority cases.
-With flexible customized kernels, it adapts well in high-paced development and test context.
+  • We propose Kernel-U-Net, a symmetric and hierarchical architecture that partitions input sequences into small patches at each layer of the network.
+  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the design pattern.
+  • The computation complexity is guaranteed in linear when employing Transformer kernels at the second or higher layers.
+  • Kernel-U-Net exceeds or meets the state-of-the-art results in most cases.
+We conclude that Kernel-U-Net stands out as a highly promising option for large-scale time series forecasting tasks.
+Its symmetric and hierarchical design provides a balance of low computational complexity and high expressiveness.
+In most scenarios, it either surpasses or is slightly below the state-of-the-art results.
+Furthermore, its adaptability in fast-paced development and testing environments is ensured by the use of flexible, customizable kernels.
+Figure 1: Comparison of the structures of different U-Net. a) classic Convolutional U-Net, b) Kernel U-Net, c) Swin Transformer U-Net
 2 Related works
@@
 As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route.
-Bidirectional Encoder Representations from Transformers (BERT)Devlin et al.
-(2019) is designed to pre-train deep bidirectional representations from the unlabeled text by joint conditioning on both left and right context in all layers.
-Generative Pre-trained Transformer 3(GPT-3) pre-trained a massive transformer-based model using 175 billion parameters Brown et al.
-(2020) and achieved strong performance on different types of downstream natural language tasks without requiring any fine-tuning.
-Later, Vision Transformers (ViTs) Dosovitskiy et al.
+Vision Transformers (ViTs) Dosovitskiy et al.
 (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al.
@@
 (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer.
-It proposes also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck.
+To improve the computation efficiency, the authors propose also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck.
 Informer Zhou et al.
 (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences.
-It proposes a ProbSparse self-attention mechanism to efficiently replace the canonical self-attention.
+The authors propose a ProbSparse self-attention mechanism to replace the canonical self-attention efficiently.
 It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments.
@@
 Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass.
+PatchTST employed PropSparse and patches, thus reducing computation complexity to $O(\frac{L}{S}log(\frac{L}{S}))$, where $S$ is the size of the patch.
 Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing.
@@
 The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
-Recently, YFormer Madhusudhanan et al.
-(2023) proposed a U-Net-like architecture that combines coarse and fine-grained features to capture long-range effects on all scale levels.
-PatchTST Nie et al.
-(2023) adopted the patches method from ViTs and proposed a time series transformer with a linear patch layer.
-It reaches state-of-the-art results on many time series forecasting benchmarks.
-The author claims Reduction in time and space complexity of learning transformer block from $O(L^{2})$ to $O((\frac{L}{S})^{2})$, where $L$ is sequence size and $S$ is stride size.
 U-Net
@@
 U-net is composed of an encoder and a decoder.
-At the encoder phase, a long sequence is reshaped into small slices and reduced by kernels step by step into a latent vector.
-At the decoder stage, the vector is developed by kernels for generating an output with the same shape as the input.
+At the encoder phase, a long sequence is reduced by a convolutional layer and a max-pooling layer gradually into a latent vector.
+At the decoder stage, the latent vector is developed by a transposed convolutional layer for generating an output with the same shape as the input.
 With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
@@
 (2016), 1D-U-Net Azar et al.
-(2022), Weighted Res-UNet Xiao et al.
-(2018) with residual structure and weighted attention mechanism, U-Net++ Zhou et al.
-(2018) with nested U-net inside each layer, UNet3+ Huang et al.
-(2020a) that allows full-scale skip connections incorporating low-level details with high-level semantics from feature maps in different scales and Swin-U-Net Cao et al.
+(2022) and Swin-U-Net Cao et al.
 (2023) that only use transformer blocks at each layer of U-Net.
-Modification on U-Net layers has been studied by Selective Kernel U-Net Byra et al.
-(2020) which merges information from different sizes of convolutional kernels and Large Kernel U-Net Jia et al.
-(2022) which combines TransMorph layers for better performance in medical image segmentation.
 In time series processing, U-time Perslev et al.
@@
 Hierarchical and hybrid model
-In time series processing, learning sequences of length $L$ requires $O(L)$ computations for recurrent models such as RNN, LSTM, GRU, and $O(L^{2})$ computations for transformer-based models.
-With ProbSparse Attention, the complexity can be reduced to $O(Llog(L))$.
-To balance the expressivity of complex modules and computational efficiency, researchers investigated hierarchical architecture and hybrid methods that merge different modules into the network.
+In time series processing, the increasing size of data degrades the performance of deep models and increases crucially the cost of learning them.
+For example, recurrent models such as RNN, LSTM, GRU have a linear complexity but suffer the gradient vanishing problem when input length increases.
+Transformer-based block captures better long dependencies but requires $O(L^{2})$ computations in general.
+To balance the expressiveness of complex models and computational efficiency, researchers investigated hybrid models that merge different modules into the network and hierarchical architectures.
 For example, authors in Du et al.
 (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al.
-(2017) applied hierarchical LSTM and GRU was used for document classification.
-In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification, authors in LogTrans Li et al.
-(2019), Informer Zhou et al.
-(2021)) and Pyraformer Liu et al.
-(2022a) studied hierarchical or pyramidal transformers for time series forecasting.
-Kernel-U-Net is a general hierarchical and symmetrical framework with low complexity in terms of parameter size.
-By replacing the convolutional kernel in classical U-Net, our model allows users to define custom kernels such as Attention, LSTM, or simple linear multiplication with a non-linear activation function for extracting features on complex time series.
-There are 3 advantages in Kernel-U-Net: 1) A small number of parameters: the total parameter of model is in order of $O(log(L)^{2})$, 2) Flexibility: the kernels can be replaced easily by arbitrary function that follows the same design mode so that non-linear kernels can be easily combined into it, 3) Computation efficiency: the computation complexity of transformer modules is reduce to $O(log(L)^{2})$ if they placed close to latent vector.
-Algorithm 1 Kernel Wrapper
-Input: kernel $f_{i}$, input segments $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
-Output: Instance of Kernel Wrapper
-1:  # Initiation of main kernel function $f$ is
-2:  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
-3:  # Holds variables for global operation
-4:  _unet_skip_output = None
-5:  _unet_skip_input = None
-6:
-7:  # Define the forward function as :
-8:  Reshape $x$ into (-1, $L_{in}$,$M_{in}$)
-9:  if _unet_skip_input is not None then
-10:     $x=x+\_unet\_skip\_input$
-11:  end if
-12:  # Apply $f$ on $x$
-13:  $y=f_{i}(x)$
-14:  _unet_skip_output = $y$
-15:  return $y$
+(2017) applied hierarchical LSTM and GRU for document classification.
+In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
+To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Duan et al.
+(2022), pyramidal structure in Pyraformer Liu et al.
+(2022a), AutoEncoders Asres et al.
+(2022), U-Net-like model in Yformer Madhusudhanan et al.
+(2023) or patches in PatchTST Nie et al.
+(2023).
+Kernel-U-Net is a symmetric and hierarchical U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences.
+Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels.
+By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data.
+Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
 3 Method
@@
 We call $X\in R^{N\times M\times L}$ the trajectory matrix which represents the sliced historical data for each instant $t\in[0,N-L-1]$.
-Time series forecasting.
-In the case of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$.
+In the context of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$.
 Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window.
@@
 Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
-Algorithm 2 Kernel U-Net Encoder
-Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
-Output: Instance of Kernel U-Net Encoder
-1:  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
-2:  # Create first layer
-[…]
-4:  # Create intermediate layers
-[…]
-6:     layers.append(KernelWrapper($f_{i},L_{i},M_{i},1,M_{i+1}$))
-7:  end for
-8:  # Create last layer
-9:  layers.append(KernelWrapper($f_{n+m},L_{n+m},M_{n+m},$ $L_{h},M_{h}$))
-10:  encoder = nn.Sequential (*layers)
 3.2 Kernel U-Net
-Kernel U-Net follows the hierarchical structure of convolutional U-Net.
-It accepts a custom kernel to replace the linear kernel inside the original convolutional U-Net.
-More precisely, it repeatedly applies custom kernels on small slices of large-scale input data.
-In this context, Kernel U-Net avoids problems in processing long sequences such as gradient vanishing in recurrent models and quadradic complexity in transformers-based models.
-In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence, and the creation of kernel wrapper, encoder, decoder, and U-net.
+Kernel U-Net(K-U-Net) is a hierarchical and symmetric U-shape architecture.
+It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels(Figure 1).
+More precisely, the K-U-Net encoder reshapes the input sequences into a large batch of small patches and repeatedly applies custom kernels on them until obtaining the latent vectors.
+Later, the K-U-Net decoder develops the batch of latent vectors into patches gradually at each layer and obtains a large batch of small patches(Figure 2).
+At last, K-U-Net reshapes the patches to get the final output.
+In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
+Figure 2: Hierarchical Partition
+Algorithm 1 Kernel Wrapper
+Input: kernel $f_{i}$, input patches $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
+Output: Instance of Kernel Wrapper
+  # Define the init function :
+  def __init(inputs)__:
+  # Initiation of main kernel function $f$
+  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
+  # Holds variables for global operation
+  _unet_skip_output = None
+  _unet_skip_input = None
+  # Define the forward function :
+  def forward(x):
+  reshape $x$ into (-1, $L_{in}$,$M_{in}$)
+  if _unet_skip_input is not None then
+           $x=x+\_unet\_skip\_input$
+  end if
+  # Apply $f$ on $x$
+  $y=f(x)$
+  _unet_skip_output = $y$
+  return $y$
 Hierarchical Partition of the input sequence.
 Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step.
-Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small slices $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all slices is equal to its original trajectory matrix:
+Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small patches $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all patches is equal to its original trajectory matrix:
 $X_{t}=Concat([[X_{t,i,j}\forall i\in[1,\frac{L}{L_{1}}]]\forall j\in[1,\frac{M}{M_{1}}]])$.
-The partitioned sequences are processed by kernels in the encoder.
-Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated segments $\hat{X}_{t+L,i,j}$.
+The partitioned sequences form a large batch of patches and will be processed by kernels in the encoder.
+Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated patches $\hat{X}_{t+L,i,j}$.
 Hierarchical processing with kernels in the encoder and decoder
@@
 At the stage of the encoder, the shape of input $X$ is $(B,L,M)$.
+Let us note $L_{1}$ and $M_{1}$ as the length and feature size of a patch.
 Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem.
-We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of slices and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$.
+We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of patches and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$.
 After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer.
-Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector.
+Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector (Algorithm 2).
 At the stage of the decoder, the operations are reversed.
@@
 At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$.
-The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$.
+The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$ (Algorithm 3)..
 The algorithm is optimized in implementation because the for loop is expensive in the forward function of a PyTorch module.
@@
 The complete algorithm is described in the next section.
+Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns.
+A kernel wrapper inherits ”nn.Module” class in Pytorch and offers an interface that executes the given kernel inside the encoder and decoder.
+It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector (Algorithm 1).
+By default, a kernel receives a batch of patches in the shape of $(-1,L_{i},M_{i})$ and outputs a new batch of patches in the shape of $(-1,L_{i+1},M_{i+1})$.
+Remark that the kernels reduce the length of patches from $L_{i}$ to $1$ in the encoder and expand the length of patches from $1$ to $L_{i}$ in the decoder.
+The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
+It reshapes the input patches and checks the output shape outside of the kernel.
+Algorithm 2 Kernel U-Net Encoder
+Input: $L_{1},M_{1}$, $[L_{2},$,...,… ,$L_{n},$ $L_{n+2},$,...,… ,$L_{n+m}]$, $[M_{2},$,...,… ,$M_{n},$ $M_{n+2},$,...,… ,$M_{n+m}]$, $[f_{1},f_{2},$,...,… ,$f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
+Output: Instance of Kernel U-Net Encoder
+  # Define the init function :
+  def __init(inputs)__:
+  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
+  # Create first layer
+[…]
+  # Create intermediate layers
+[…]
+           layers.append(KernelWrapper($f_{i},$ $L_{i},$ $M_{i},$ $1,$
+           $M_{i+1}$))
+  end for
+  # Create last layer
+  layers.append(KernelWrapper($f_{n+m},L_{n+m},$
+  $M_{n+m},$ $L_{h},M_{h}$))
+  encoder = nn.Sequential (*layers)
+Creation of Kernel U-Net
+Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of patch $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$ Remark that the hidden dimension of the model is the size of intermediate output vectors within layers.
+It corresponds to channel size in a convolutional network.
+This dimension can be augmented for a larger passage of information.
+We describe the creation of the K-U-Net encoder in Algorithm 2 and leave that of the decoder (Algorithm 3) and K-U-Net (Algorithm 4) in the appendix.
+3.3 Custom Kernels
+Linear kernel
+The linear kernel is a simple matrix multiplication (Figure 3). Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
+$Z_{i}=f_{i}(X)=X*w+b$
+, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector.
+Remark that the number of parameters of $w$ in a linear kernel is $L_{i}*1*ch_{in}*ch_{out}$ which is equivalent to that of a 1D convolutional layer whose kernel size is $L_{i}*1$.
+The multi-layer linear kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
+$Z_{i}=f_{i}(X)=Tanh(X*w_{1}+b_{1})*w_{2}+b_{2}$
+, where $Z_{i}$ is output of kernel, $w_{1}\in R^{L_{i}M_{i}\times L^{\prime}M^{\prime}}$ and $w_{1}\in R^{L^{\prime}M^{\prime}\times L_{i+1}M_{i+1}}$ are weight matrix , $b_{1}\in R^{L^{\prime}M^{\prime}}$ and $b_{2}\in R^{L_{i+1}M_{i+1}}$ are bias vectors, and $L^{\prime}=\frac{1}{2}(L_{i}+L_{i+1}),M^{\prime}=\frac{1}{2}(M_{i}+M_{i+1})$.
+Transformer kernel
+The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with Relu activation.
+The transformer kernel in this work follows the classical structure.
+LSTM kernel
+The LSTM kernel contains a classic LSTM cell and a linear layer for skip connection.
+The hidden states of all time steps are flattened and used for a linear combination in the next.
+The detailed structure of the Transformer kernel and the LSTM kernel is discussed in the appendix.
+Figure 3: Structure of Linear Kernel. a) Linear kernel, b) Multi-layer Linear Kernel with Tanh activation.
+Table 1: Multivariate time series forecasting results with Kernel U-Net.
+The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
+We note the best results in bold and the second best results in underlined.
+Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1
+96 0.370 0.393 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
+192 0.404 0.414 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
+336 0.420 0.430 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
+720 0.438 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
+ETTh2
+96 0.271 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
+192 0.332 0.377 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931 1.593 1.021
+336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
+720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
+ETTm1
+96 0.286 0.342 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
+192 0.330 0.363 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
+336 0.360 0.384 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
+720 0.405 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
+ETTm2
+96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
+192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
+336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
+720 0.343 0.379 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
+Electricity
+96 0.129 0.226 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
+192 0.147 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
+336 0.163 0.261 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
+720 0.197 0.292 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
+Traffic
+96 0.358 0.253 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
+192 0.373 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
+336 0.390 0.271 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
+720 0.430 0.292 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
+Weather
+96 0.142 0.195 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
+192 0.188 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
+336 0.241 0.284 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
+720 0.310 0.333 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
+3.4 Complexity Analysis
+As the K-U-Net is symmetric, we study the complexity of the encoder.
+Let us suppose that a kernel is receiving a sequence of length $L$ where $L=\prod_{1}^{n}L_{i}$ and all $L_{i}$ are equal.
+As the patch size in the first layer of Kernel U-Net encoder is $L_{1}$, a kernel module will process $\frac{L}{L_{1}}$ patches of size $L_{1}$.
+Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}*f(L_{1}))$ where $f(L_{1})$ is the complexity inside the kernel in the function of the patch size $L_{1}$ and $n$ is the index of the layer.
+In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}*L_{1})$= $O(L)$.
+In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}*L_{1}^{2})$= $O(L*L_{1})$.
+Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(L*log(L))$.
+Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}^{2}}*L_{1}^{2})$= $O(L)$.
+Following the same demonstration, the complexity of using multi-layer linear and LSTM kernels starting from the second layer is also bounded by $O(L)$.
+Figure 4: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on ETTh1 dataset.
+4 Experiments and Results
+Datasets.
+We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^1] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Traffic [^3] http://pems.dot.ca.gov and Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams
+20112014
+. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
+Figure 5: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on Weather dataset.
+Here, we followed the experiment setting in Zeng et al.
+(2023) and partitioned the data into $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset.
+The data is split into $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
+Baselines and Experimental Settings.
+We follow the experiment setting in NLinear Zeng et al.
+(2023) and take $L=\{336,720\}$ step historical data as input then forecast $T\in\{96,192,336,720\}$ step value in the future.
+We replace the last value normalization with mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalizationUlyanov et al.
+(2017) for the Traffic dataset.
+We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as mentioned in Wu et al.
+(2021).
+We include recent methods: PatchTSTNie et al.
+(2023), NLinear, DLinearZeng et al.
+(2023), FEDformer Zhou et al.
+(2022), Autoformer Wu et al.
+(2021), Informer Zhou et al.
+(2021), LogTrans Li et al.
+(2019) YformerMadhusudhanan et al.
+(2023).
+We merge the result reported in Nie et al.
+(2023) and Zeng et al.
+(2023) for taking their best in a supervised setting and execute Yformer in our environment with default parameters in its Github [^5] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
+Figure 6: Computation time of K-U-Net, PatchTST and Yformer.
+Model Variants Search.
+We propose 4 kernels for experiments with K-U-Net on 7 datasets.
+By replacing a linear kernel at different layers with other types of kernels for searching for variants that adapt the dataset.
+The variants are noted as ”K-U-Net $<$kernel$>$_$<$replacing_layer$>$ ($<$lookback_window$>$)”.
+For example, K-U-Net Linear_$\_0000(720)$ means that the model is made of linear kernels at all layers and the lookback window size is 720, K-U-Net Transf_$\_0100(720)$ means that a transformer kernel replaces the linear kernel at the second layer, K-U-Net Linear_$\_0110(720)$ means that multilayer linear kernels replace the kernels at the second and third layers.
+We compose 16 variants with Multi-layer linear kernels and 7 variants with Transformer and LSTM kernels.
+To enumerate all variants that achieve at least once the best result, we report their performance by the average of the top 5 minimum running MSE(Top5MMSE) values on the validation and test set.
+The search results are reported with a relative score(RS) to the minimum Top5-MMSE:
+$RS(m)=Min_{t\in T}\frac{Top5MMSE(m,t)}{Min_{\{m\in M,l\in L\}}Top5MMSE(m,t,l)}$
+, where $m$ is a model in the models set $M$, $t$ is a forecasting horizon in $T$, $l$ is a look back window in $L$.
+Relative score notes the best-performed model with 1 and thus helps to identify the high-potential candidates for further fine-tuning examinations.
+We observe in Figure 4 that the best variant for ETTh1 dataset is K-U-Net Linear_$\_0010$.
+Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting.
+Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are K-U-Net Linear_$\_0110$, K-U-Net LSTM_$\_0100$ and K-U-Net Transf_$\_0100$.
+Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
+Among candidates in the search phase, we empirically choose 3 variants for fine-tuning experiments with 5 runs.
+The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in univariate settings and multivariate settings in most cases.
+Multivariate time series forecasting result.
+We remark that our model improves the MSE performance around $72.92\%$ compared with Yformer and $0.99\%$ compared to PatchTST and NLinear in the multivariate setting (Table 1).
+It is worth noting that K-U-Net achieves similar results on the Electricity dataset with a variant based on Multilayer linear kernels.
+Univariate time series forecasting result.
+Additionally, we remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3).
+It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ compared to PatchTST.
+Figure 7: GPU consumation of K-U-Net, PatchTST and Yformer.
+Computation Efficiency
+We examined the computation efficiency of 4 variants of K-U-Net, PatchTST, and Yformer.
+We execute the models on the ETTh1 and ETTm1 datasets for 10 epochs and measure the average execution time per epoch and the GPU consumption during the training.
+The hidden dimension is set to 128 for all models.
+For fair comparison, PatchTST and Yformer contain 2 layers of Transformer block which equals to K-U-Net_Transf_$\_0100$.
+All experiments are executed on a Tesla V100 GPU in a Google Colab environment.
+Comparing to the PatchTST, K-U-Net Linear_$\_0000$ saves $85.20\%$ and $83.07\%$ memory (Figure 7) and saves $81.82\%$ and $80.55\%$ computation time (Figure 6) on ETTh1 and ETTm1 datasets, K-U-Net_Transf_$\_0100$ saves $55.15\%$ and $50.00\%$ memory and saves $28.47\%$ and $26.88\%$ computation time respectively.
+5 Conclusion
+In this paper, we propose Kernel-U-Net, a highly potential candidate for large-scale time series forecasting tasks.
+It provides convenience for composing particular models with custom kernels, thereby it adapts well to particular datasets.
+As an efficient architecture, it accelerates the procedure of searching for appropriate variants.
+Kernel-U-Net either exceeds or meets the state-of-the-art results in most cases.
+In the future, we hope to develop more kernels and hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.
+Appendix A Appendix
+A.1 Complement Algorithms
+The complete algorithm of Kernel U-Net contains creation of Encoder, Decoder (Algorithm 3), Wrapper, and Kernel U-Net(Algorithm 4).
 Algorithm 3 Kernel U-Net Decoder
@@
 Output: Instance of Kernel U-Net Decoder
-1:  # Create a list of multiples
+  # Define the init function :
+  def __init(inputs)__:
+  # Create a list of multiples
 […]
-3:  # Create first layer
-4:  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$ $L_{n+m},$ $M_{n+m}$)]
-5:  # Create intermediate layers
-6:  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
-[…]
-8:  end for
-9:  # Create last layer
-10:  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
-11:  decoder = nn.Sequential (*layers)
-Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns.
-A kernel wrapper is an interface that executes the given kernel inside the encoder and decoder.
-It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector.
-By default, a kernel reduces the length of a segment from $L_{i}$ to $1$ in the encoder and develops the length of a segment from $1$ to $L_{i}$ in the decoder.
-The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
+  # Create first layer
+  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$
+  $L_{n+m},$ $M_{n+m}$)]
+  # Create intermediate layers
+  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
+           layers.append(KernelWrapper($f_{i},$ $1,M_{i+1}$
+           $,L_{i},M_{i})$)
+  end for
+  # Create last layer
+  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
+  decoder = nn.Sequential (*layers)
 Algorithm 4 Kernel U-Net
@@
 Output: Instance of Kernel U-Net
-1:  # Initiation of encoder and decoder in K-U-N.
-2:  encoder = KernelUNetEncoder(inputs)
-3:  decoder = KernelUNetDecoder(inputs)
-4:
-5:  # Define the forward function as
-6:  z = encoder(x)
-7:  for  $i$ in $[1,...,n]$ do
-8:     output = encoder.layers[i]._unet_skip_output
-9:     decoder.layers[$n-i+1$]. _unet_skip_input = output
-10:  end for
-11:  y= decoder(z)
-12:  return y
-Creation of Kernel U-Net
-Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of segment-unit $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$.
-Remark that the hidden dimension of the model is the size of intermediate output vectors within layers.
-It corresponds to channel size in a convolutional network.
-This dimension can be augmented for a larger passage of information.
-The kernel wrapper is described in algorithm 1.
-The creation of Kernel U-Net Encoder, Decoder, and complete U-Net are described in algorithms 2, 3, 4.
-Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.369 0.404 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713
-192 0.397 0.426 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792
-336 0.429 0.428 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809
-720 0.433 0.45 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865
-ETTh2
-96 0.271 0.336 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525
-192 0.332 0.376 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931
-336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835
-720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625
-ETTm1
-96 0.285 0.337 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571
-192 0.327 0.367 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669
-336 0.357 0.392 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871
-720 0.408 0.422 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823
-ETTm2
-96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453
-192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563
-336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887
-720 0.348 0.378 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338
-Electricity
-96 0.129 0.225 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368
-192 0.148 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386
-336 0.168 0.263 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394
-720 0.205 0.299 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439
-Traffic
-96 0.360 0.252 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391
-192 0.377 0.263 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379
-336 0.389 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42
-720 0.427 0.290 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472
-Weather
-96 0.146 0.200 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384
-192 0.192 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544
-336 0.241 0.287 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523
-720 0.310 0.340 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741
-Table 1: Multivariate time series forecasting results with Kernel U-Net.
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold and the second best results in underlined.
-3.3 Custom Kernels
-Linear kernel
-The linear kernel is a simple matrix multiplication. Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
-$Z_{i}=f_{i}(X)=X*w+b$
-, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector.
-Remark that linear Kernel U-Net is an equivalent version to convolutional U-Net whose stride size is $L_{i}$.
-We also test a non-linear kernel that takes a non-linear activation function such as Tanh in the intermediate hidden layers.
+  # Define the init function :
+  def __init(inputs)__:
+  # Initiation of encoder and decoder in K-U-N.
+  encoder = KernelUNetEncoder(inputs)
+  decoder = KernelUNetDecoder(inputs)
+  # Define the forward function :
+  def forward(x):
+  reshape x to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
+  transpose x to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
+  reshape x to $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$
+  # Apply encoder on x
+  z = encoder(x)
+  for  $i$ in $[1,...,n]$ do
+           output = encoder.layers[i]._unet_skip_output
+           decoder.layers[$n-i+1$]._unet_skip_input
+           = output
+  end for
+  # Apply decoder on z
+  y= decoder(z)
+  reshape y to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
+  transpose y to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
+  reshape y to $(B,L,M)$
+  return y
+A.2 LSTM and Transformer kernels
 Transformer kernel
-The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.Wen et al. (2023)
+The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.
+We follow the description in Wen et al.
+[2023].
 Positional encoding (PE).
@@
 By adding positional information to the input, the transformer can consider the same patterns that occur in different time steps as different patterns.
+Figure 8: Structure of Kernels a) Linear Kernel, b) Multi layer linear Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
 Multiple head attentions.
@@
 , where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values.
-Remark that in the context of self-attention, $Q=K=V=inputdata$.
 The multiple attention block contains repeated heads and can be given by
@@
 , where $h_{i}=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}).$
+The transformer kernel in this work (Figure 8) follows the classical structure with 2 additional linear layers.
+The first linear layer before the Transformer block of size $M_{in}\times M_{out}$ reduces the dimension.
+The last linear layer is a matrix of size $L_{in}*M_{out}\times L_{out}*M_{out}$ that reduces the input length.
 LSTM kernel
@@
 The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state.
-The formulation of LSTM is as follows:
+The formulation of the LSTM cell is as follows:
 $f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$
@@
 $h_{t}=o_{t}tanh(C_{t})$
-Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer    LogTrans
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.052 0.176 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468
-192 0.067 0.202 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409
-336 0.074 0.22 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546
-720 0.074 0.223 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629
-ETTh2
-96 0.114 0.262 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379
-192 0.149 0.304 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429
-336 0.158 0.319 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437
-720 0.178 0.34 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387
+In our experiments, we add linear layers (Figure 8) that take the flattened output of all hidden states of the LSTM cell for linear combination.
+$Z=W*Flatten([h_{1},...,h_{t}])+b$
+, where $W\in R^{L_{i}*d_{h}}$, $b\in R^{L_{i+1}*M_{i+1}}$ and $d_{h}$ is the dimension of hidden state and $d_{h}=M_{i}$ in general.
+The residual skip operation aside from the LSTM cell contains a linear layer equivalent to a simple linear kernel discussed in the previous sections.
+A.3 Dataset and Experiment
+Dataset Statistics
+We use 7 popular multivariate datasets for experiments.
+We create subset Weather(S), Electricity(S), and Traffic(S) from top-left to bottom-right for quick search of K-U-Net variants.
+The statistic is in noted in Table 2.
+We refer the description of datasets in Nie et al.
+[2023]:
+  • Weather dataset collects 21 meteorological indicators in Germany, such as humidity and air temperature.
+  • Traffic dataset records the road occupancy rates from different sensors on San Francisco freeways.
+  • Electricity is a dataset that describes 321 customers’ hourly electricity consumption.
+  • ETT (Electricity Transformer Temperature) datasets are collected from two different electric transformers labeled with 1 and 2, and each of them contains 2 different resolutions (15 minutes and 1 hour) denoted with m and h.
+Table 2: Statistics of popular datasets for experiments.
+Datasets ETTh1 ETTh2
 ETTm1
-96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171
-192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317
-336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459
-720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579
-ETTm2
-96 0.061 0.181 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208
-192 0.085 0.22 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275
-336 0.104 0.25 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302
-720 0.149 0.303 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321
+ETTm2 Weather
+Traffic
+Electricity Weather(S) Traffic(S) Electricity(S)
+Features 7 7
+7
+7 21
+862
+321 21 (100%) 21 (2.5%) 22 (7%)
+Timesteps 17420 17420
+69680
+69680 52696
+17544
+26304 10539 (20%) 12280 (70%) 10521 (40%)
 Table 3: Univariate time series forecasting results with Kernel U-Net.
 The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold and the second best results in underlined.
-The LSTM Kernel used in our experiment contains an LSTM cell and a linear layer that takes all hidden states of a sequence as input and gives a vector as output.
-The experimental result is shown in the section 4
-4 Experiments and Results
-Datasets.
-We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Traffic and Electricity.
-These datasets have been benchmarked and publicly available onWu et al.
-(2021) and the description of the dataset is available in Zeng et al.
-(2023)).
-Here, we followed the experiment setting in Zeng et al.
-(2023) and partitioned the data into 12, 4, and 4 months for training, validation, and testing respectively for the ETT dataset.
-The data is split into [0.7,0.1,0.2] for training, validation, and testing for dataset Weather, Traffic, and Electricity.
-Baselines and Experimental Settings.
-In the time series forecasting task, we follow the experiment setting in NLinear Zeng et al.
-(2023), which take $L={336,720}$ step historical data as input and forecast $T\in{96,192,336,720}$ step value in the future.
-We replace the last value normalization with mean value normalization for univariate forecasting in the ETT dataset.
-We include five recent methods: PatchTSTNie et al.
-(2023), NLinear, DLinearZeng et al.
-(2023), FEDformer Zhou et al.
-(2022), Autoformer Wu et al.
-(2021), Informer Zhou et al.
-(2021)and LogTrans Li et al.
-(2019).
-We merge the result reported in Nie et al.
-(2023) and Zeng et al.
-(2023) for taking their best in a supervised setting.
-Following previous works Wu et al.
-(2021), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the core metrics to compare performance for Forecasting problems.
-We take a look back at Windows 336 and 720 in experiments.
-The list of multiples is respectively [3, 4, 4, 5] and [4, 4, 7] and the segment-unit input length is 3.
-The input dimension is 1 as we follow the channel-independent setting in Zeng et al.
-(2023).
-The hidden dimensions are 128 or 256 for multivariate tasks and 5 or 30 for univariate tasks depending on of dataset.
-We use the Random Erase Zhong et al.
-(2020) from image processing for data augmentation.
+We note the best results in bold and the second best in underlined.
+Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer LogTrans Yformer
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1
+96 0.053 0.177 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468 0.211 0.379
+192 0.069 0.205 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409 0.228 0.403
+336 0.076 0.221 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546 0.179 0.355
+720 0.077 0.224 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629 0.260 0.444
+ETTh2
+96 0.114 0.267 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379 0.240 0.398
+192 0.146 0.306 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429 0.270 0.429
+336 0.157 0.321 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437 0.263 0.422
+720 0.182 0.347 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387 0.264 0.423
+ETTm1
+96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171 0.211 0.393
+192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317 0.262 0.42
+336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459 0.424 0.559
+720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579 0.413 0.552
+ETTm2
+96 0.062 0.182 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208 0.128 0.275
+192 0.089 0.224 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275 0.155 0.308
+336 0.114 0.259 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302 0.208 0.358
+720 0.156 0.312 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321 0.261 0.405
+Table 4: Comparison of 16 variants of Linear K-U-Net in baseline setting and without using Skip connection(-Skip), Random Erase(-RE), Chanel Independence(-CI).
+The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
+We note the best results in bold and the second best results in underlined for minimum score(Min) and average score(Mean).
+Method Min (Base) Min (-Skip) Min (-RE) Min (-CI) Mean (Base) Mean (-Skip) Mean (-RE) Mean (-CI)
+Metric MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE MSE Top5-
+MMSE
+ETTh1
+96 0.368 0.368 0.375 0.375 0.370 0.369 0.459 0.451 0.376 0.374 0.393 0.385 0.377 0.373 0.488 0.483
+192 0.404 0.402 0.412 0.413 0.402 0.402 0.506 0.495 0.409 0.407 0.423 0.423 0.411 0.407 0.551 0.538
+336 0.423 0.421 0.443 0.441 0.427 0.422 0.535 0.519 0.443 0.430 0.489 0.455 0.442 0.431 0.581 0.549
+720 0.457 0.439 0.461 0.453 0.461 0.438 0.629 0.576 0.492 0.472 0.531 0.515 0.488 0.470 0.690 0.598
+ETTh2
+96 0.274 0.272 0.273 0.273 0.274 0.273 0.423 0.397 0.302 0.287 0.299 0.290 0.300 0.286 0.458 0.415
+192 0.336 0.353 0.337 0.355 0.336 0.354 0.414 0.442 0.370 0.375 0.376 0.382 0.372 0.374 0.445 0.474
+336 0.361 0.361 0.364 0.364 0.362 0.361 0.412 0.422 0.402 0.384 0.404 0.392 0.399 0.384 0.428 0.444
+720 0.398 0.394 0.397 0.397 0.395 0.394 0.453 0.456 0.451 0.418 0.459 0.436 0.444 0.420 0.468 0.519
+ETTm1
+96 0.288 0.284 0.289 0.290 0.288 0.284 0.321 0.318 0.296 0.292 0.303 0.298 0.296 0.291 0.346 0.333
+192 0.324 0.327 0.325 0.324 0.326 0.324 0.359 0.358 0.335 0.331 0.341 0.336 0.334 0.331 0.383 0.373
+336 0.359 0.358 0.353 0.354 0.359 0.357 0.391 0.391 0.366 0.363 0.374 0.368 0.366 0.362 0.418 0.406
+720 0.408 0.404 0.403 0.403 0.409 0.403 0.452 0.447 0.418 0.411 0.419 0.413 0.418 0.410 0.478 0.467
+ETTm2
+96 0.162 0.162 0.162 0.161 0.163 0.161 0.213 0.204 0.168 0.165 0.170 0.165 0.168 0.164 0.228 0.210
+192 0.218 0.217 0.216 0.215 0.217 0.216 0.298 0.274 0.228 0.223 0.229 0.224 0.228 0.222 0.318 0.286
+336 0.270 0.268 0.268 0.267 0.265 0.265 0.372 0.345 0.283 0.277 0.281 0.277 0.284 0.272 0.394 0.356
+720 0.342 0.349 0.343 0.347 0.343 0.347 0.410 0.408 0.354 0.361 0.351 0.361 0.364 0.356 0.434 0.421
+Average decrease +0.42% +0.77% +0.03% -0.20% +24.19% +21.45% +2.24% +2.30% +0.03% -0.32% +25.92% +23.72%
+Experiment details
+We use a 4-layer K-U-Net for experiments.
+The list of multiples are respectively [4, 4, 3, 7] and [4, 6, 6, 5] for look-back windows $L$ and forecasting horizons $T$ in {336, 720}.
+Remark that $T$ is fixed to be 336 for shorter forecasting cases.
+The bottom patch length is 4 and its width is 1 as we follow the Channel-Independent (CI) setting in Zeng et al.
+[2023].
+In case of not using CI, the patch width equals to feature size.
+The hidden dimensions are 128 for multivariate and univariate tasks in general, except that they are 256 on the Traffic dataset.
 The learning rate is selected in $[0.00005,0.001]$ depending on the dataset.
-The training epoch is 300 and the patience is 30 in general.
-Instance Normalization and Mean Normalization Instance normalization Ulyanov et al.
-(2017) Kim et al.
-(2021) contains a step of substruction with the average value of the input sequence and a division of its standard deviation.
+The training epoch is 50 and the patience of early stopping is 10 in general.
+We apply weighted early stopping where the triggering MSE (TMSE) is computed with $TMSE_{e+1}=\tau*TMSE_{e}+(1-\tau)*VMSE_{e}$ , where $\tau$ is a hyper-parameter to be set, $e$ is the index of the episode, and VMSE is the running MSE on a validation set.
+The parameter $\tau$ sets the sensitivity of early stopping and is empirically selected.
+In general, it is set to be 0.9 in the variants searching stage and 0.5 in the fine-tuning stage.
+The MSE loss function is defined with $:\mathcal{L}=\mathbb{E}_{\boldsymbol{x}}\frac{1}{B*M}\sum_{i=1}^{B*M}\left\|\hat{\boldsymbol{x}}_{L+1:L+T}^{(i)}-\boldsymbol{x}_{L+1:L+T}^{(i)}\right\|_{2}^{2}$ , where $B$ is the batch size and $M$ is the multiple of feature size unit of time series.
+A.4 Search of model variants
+We searched 30 variants of K-U-Net across Linear_[$0$,…,$1$] and {LSTM, Transformer}_[0100,…,$1$] for the ETT, Weather datasets and 3 subsets.
+The lookback window is typically 720, and we additionally tested a version of 336 on the ETT dataset.
+The result is reported with the average of running Top5 Minimum MSE (Top5MMSE) on the validation and test set in a single run.
+This measures the extreme performance of a variant and its potential in the fine-tuning stage.
+We use relative score (RS) to locate best-performed variants (i.g.
+$RS=1$ means that it’s the best among all candidates).
+We plot the RS of variants of 5 datasets in Figures 9 and Figures 10.
+Based on the search result, We note that Linear and Multilayer linear kernels are more adapted in the ETT dataset than our Transformer and LSTM kernels.
+We remark that K-U-Net Linear_1000 is outstanding in the Electricity (S) dataset and thus selected for full examination.
+A.5 Result of selected variants
+In Table 5 6 7, we note selected models that are examined prudently with 5 runs in a fine-tuned context and measures with average MSE and MAE.
+We include the Table 9 showing that K-U-Net has a small standard deviation error across 5 random runs.
+A.6 Result of univariate time series forecasting
+The experiment of univariate time series forecasting task predicts the target feature is the ”oil temperature” on ETT datasets.
+We remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3).
+It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ in the Univariate setting compared to PatchTST.
+Appendix B Ablation Study
+For studying the roles of skip connection in K-U-Net, Random Erase, and Chanel Independence, we use the Minimum Top5 average MSE (Min Top5MMSE) to measure the extreme performance and the Mean Top5MMSE to measure the average performance.
+For instance Normalisation, we only show the comparison with Mean Normalization with the Traffic dataset because the improvement is evident.
+Table 5: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm.
+The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
+We note the best results in bold.
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_0000
+(720) K-U-Net
+Linear_1100
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1
+96 0.372 0.393 0.377 0.401 0.370 0.407 0.370 0.393 0.370 0.400
+192 0.405 0.414 0.413 0.424 0.404 0.433 0.404 0.414 0.413 0.429
+336 0.431 0.430 0.444 0.444 0.420 0.445 0.420 0.430 0.422 0.440
+720 0.438 0.454 0.457 0.471 0.482 0.490 0.438 0.454 0.447 0.468
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_0000
+(720) K-U-Net
+Linear_0001
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh2
+96 0.273 0.335 0.271 0.336 0.286 0.342 0.271 0.335 0.274 0.337
+192 0.336 0.377 0.332 0.377 0.354 0.384 0.332 0.377 0.339 0.379
+336 0.360 0.401 0.357 0.403 0.388 0.413 0.357 0.401 0.329 0.384
+720 0.395 0.435 0.394 0.442 0.420 0.447 0.394 0.435 0.379 0.422
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_0000
+(720) K-U-Net
+Linear_1000
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm1
+96 0.301 0.343 0.310 0.351 0.286 0.342 0.286 0.342 0.290 0.342
+192 0.336 0.363 0.338 0.368 0.330 0.371 0.330 0.363 0.332 0.369
+336 0.371 0.384 0.367 0.385 0.360 0.393 0.360 0.384 0.366 0.392
+720 0.427 0.416 0.417 0.412 0.405 0.420 0.405 0.412 0.416 0.420
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_0000
+(720) K-U-Net
+Linear_0110
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm2
+96 0.164 0.253 0.161 0.252 0.165 0.256 0.161 0.252 0.165 0.255
+192 0.219 0.290 0.216 0.290 0.225 0.296 0.216 0.290 0.220 0.292
+336 0.273 0.326 0.268 0.326 0.275 0.332 0.268 0.326 0.274 0.329
+720 0.367 0.383 0.350 0.379 0.343 0.380 0.343 0.379 0.362 0.385
+Table 6: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets Electricity, Weather and Traffic.
+The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
+We note the best results in bold.
+Method K-U-Net
+Linear_0000
+(720) K-U-Net
+Linear_1110
+(720) K-U-Net
+Linear_1000
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+Electricity
+96 0.138 0.237 0.134 0.232 0.129 0.226 0.129 0.226 0.129 0.222
+192 0.153 0.250 0.152 0.249 0.147 0.244 0.147 0.244 0.147 0.240
+336 0.171 0.269 0.167 0.264 0.163 0.261 0.163 0.261 0.163 0.259
+720 0.213 0.305 0.203 0.300 0.197 0.292 0.197 0.292 0.197 0.290
+Method K-U-Net
+Linear_0000
+(720) K-U-Net
+Transf_0001
+(720) K-U-Net
+LSTM_0001
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+Traffic
+96 0.412 0.280 0.399 0.262 0.358 0.253 0.358 0.253 0.360 0.249
+192 0.425 0.287 0.415 0.278 0.373 0.262 0.373 0.262 0.379 0.250
+336 0.438 0.294 0.424 0.295 0.390 0.271 0.390 0.271 0.392 0.264
+720 0.458 0.314 0.465 0.309 0.430 0.292 0.430 0.292 0.432 0.286
+Method K-U-Net
+Linear_0110
+(720) K-U-Net
+Transf_0001
+(720) K-U-Net
+LSTM_0100
+(720) MIN PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+Weather
+96 0.146 0.201 0.146 0.200 0.142 0.195 0.142 0.195 0.149 0.198
+192 0.190 0.245 0.192 0.246 0.188 0.244 0.188 0.244 0.194 0.241
+336 0.241 0.284 0.242 0.287 0.242 0.286 0.241 0.284 0.245 0.282
+720 0.310 0.333 0.310 0.336 0.316 0.341 0.310 0.333 0.314 0.334
+B.1 Instance Normalisation and Mean Normalisation
+Instance normalization contains a step of substruction with the average value of the input sequence and a division of its standard deviation.
 It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model.
-Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises in the small dataset.
-We followed the mean normalization as the preprocessing step and compared it with instance normalization for selected models.
-Model Variants.
-We propose 4 kernels and 5 versions of kernel U-Net for experiments. The kernels tested are Linear, Linear with 1 hidden layer and non-linear activation, Transformer, and LSTM.
-  • Linear U-Net is a baseline method that contains a list of linear kernels.
-  • Linear-1-Hidden U-Net replaces 2 kernels close to latent vector to kernels with hidden layer and nonlinear activation.
-  • Linear-5-Hidden U-Net replaces all linear kernels with non-linear kernels.
-  • Transformer U-Net replaces 2 kernels close to latent vector to transformer kernels and keeps the rest as linear kernels.
-  • LSTM U-Net replaces 2 kernels close to latent vector to LSTM kernels and keeps the rest as linear kernels.
-The result shows that the performance of Kernel U-Net reaches or outperforms the state-of-art methods in univariate setting (Table 2) and reaches similar or better results in multivariate setting (Table 1).
-5 Conclusion
-In this paper, we propose Kernel U-Net, a symmetric and hierarchical framework that can effectively process long temporal dependencies with $O(log(L)^{2})$ number of parameters.
-Specifically, Kernel U-Net cut input sequences into small slices at each layer in the network for applying customized kernels.
-In the future, we hope to develop more kernels that adapt well to particular datasets.
-We also hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.
+Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises into the dataset.
+We employed the mean normalization as the preprocessing step for most experiments except those with instance normalization on traffic datasets.
+The comparison in Table 8 shows that Instance Normalization improves the MSE by 3.89% for a K-U-Net LSTM_0001 on traffic datasets.
+Table 7: Fine-tuned univariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm.
+The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
+We note the best results in bold.
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_1111
+(336) MIN PatchTST Nlinear
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1
+96 0.053 0.177 0.054 0.182 0.053 0.177 0.055 0.179 0.053 0.177
+192 0.069 0.205 0.079 0.219 0.069 0.205 0.071 0.215 0.069 0.204
+336 0.076 0.222 0.076 0.221 0.076 0.221 0.076 0.22 0.081 0.226
+720 0.077 0.224 0.078 0.226 0.077 0.224 0.087 0.232 0.08 0.226
+Method K-U-Net
+Linear_0110
+(336) K-U-Net
+Linear_1100
+(336) MIN PatchTST Nlinear
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh2
+96 0.114 0.267 0.121 0.272 0.114 0.267 0.129 0.282 0.129 0.278
+192 0.146 0.306 0.161 0.317 0.146 0.306 0.168 0.328 0.169 0.324
+336 0.157 0.321 0.178 0.340 0.157 0.321 0.171 0.336 0.194 0.355
+720 0.192 0.355 0.182 0.347 0.182 0.347 0.223 0.38 0.225 0.381
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_0000
+(720) MIN PatchTST Nlinear
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm1
+96 0.026 0.121 0.026 0.122 0.026 0.121 0.026 0.121 0.026 0.122
+192 0.038 0.149 0.039 0.150 0.038 0.149 0.039 0.15 0.039 0.149
+336 0.051 0.173 0.051 0.174 0.051 0.173 0.053 0.173 0.052 0.172
+720 0.072 0.206 0.068 0.203 0.068 0.203 0.073 0.206 0.073 0.207
+Method K-U-Net
+Linear_0000
+(336) K-U-Net
+Linear_1000
+(720) MIN PatchTST Nlinear
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTm2
+96 0.063 0.182 0.062 0.183 0.062 0.182 0.065 0.187 0.063 0.182
+192 0.090 0.224 0.089 0.227 0.089 0.224 0.093 0.231 0.09 0.223
+336 0.116 0.259 0.114 0.263 0.114 0.259 0.121 0.266 0.117 0.259
+720 0.170 0.319 0.156 0.312 0.156 0.312 0.171 0.322 0.17 0.318
+Table 8: Comparison of Mean Normalization(+MN) and Instance Normalization (+IN) on Traffic dataset. We note the best results in bold..
+Methods K-U-Net (+MN)
+LSTM_0001
+(720) K-U-Net (+IN)
+LSTM_0001
+Metric MSE MAE MSE MAE
+Traffic
+96 0.375 0.264 0.360 0.252
+192 0.393 0.275 0.377 0.263
+336 0.406 0.280 0.389 0.270
+720 0.442 0.296 0.427 0.290
+Table 9: MSE and standard deviation of K-U-Net across 5 random runs.
+Method K-U-Net
+LSTM_0001
+(720) Method K-U-Net
+LSTM_0100
+(720) Method K-U-Net
+Linear_1000
+Metric MSE Metric MSE Metric MSE
+Traffic
+96 0.35761 ±0.00042
+Weather
+96 0.14200 ±0.00016
+ETTm1
+96 0.28600 ±0.00056
+192 0.37266 ±0.00056 192 0.18782 ±0.00078 192 0.32996 ±0.00204
+336 0.39030 ±0.00135 336 0.24164 ±0.00139 336 0.35979 ±0.00092
+720 0.42970 ±0.00078 720 0.31646 ±0.00198 720 0.40539 ±0.00128
+B.2 Skip Connection,
+Skip Connection in the U-Net is used for passing information from the encoder to the same level on its symmetric decoder.
+As the size of the hidden vector is reduced during passage in the encoder, skip connection helps keep the low-level features and allows the decoder to mix with high-level information.
+We conduct experiments to understand the role of skip connection in the K-U-Net.
+The result in Table 4 shows that removing the skip connection degrades the Min Top5MMSE score for $0.77\%$ on average and especially $3.19\%$ on the ETTh1 dataset.
+Additionally, it degrades $2.30\%$ Mean Top5MMSE on average and especially $5.42\%$ on the ETTh1 dataset.
+B.3 Comparison of Chanel Independent and Chanel Mixing
+Chanel Independent was an inherited choice from Zeng et al.
+[2023].
+It avoids introducing noises that degrade the performance across different features.
+In this experiment, we compare the channel-independent(CI) strategy and the channel-mixing strategy.
+In K-U-Net, setting input and output dimension to 1 implies a channel-independent strategy, while setting to feather size (7 for the ETT dataset) means a channel-mixing strategy.
+As We observe in Table 4, removing CI degrades the Min Top MMSE score for $21.45\%$ and the Min Top5MMSE score for $23.72\%$.
+B.4 Comparison of Random Erasing
+Random Erasing augments the dataset by masking partially the input.
+In this experiment, we studied the role of random erasing in time series forecasting settings.
+We observe in Table 4 that removing Random Erasing degrades the Min Top5MMSE score for $0.2\%~{}0.32\%$, which is not as evident as the Skip Connection and Channel Independent strategy.
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Weather
+(f) Weather(S)
+(g) Electricity(S)
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Weather
+(f) Weather(S)
+(g) Electricity(S)
+(h) Traffic(S)
+Figure 9: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on validation and test set.
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Weather
+(f) Weather(S)
+(g) Electricity(S)
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Weather
+(f) Weather(S)
+(g) Electricity(S)
+(h) Traffic(S)
+Figure 10: Search result of K-U-Net with Multi-layer Linear kernels on validation and test set.

### v3

@@
-Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting
+Kernel-U-Net: Multivariate Time Series Forecasting using Custom Kernels
 Abstract
@@
 Transformer-based U-Net architectures, despite their success in medical image segmentation, have limitations in both expressiveness and computation efficiency in time series forecasting as evidenced in YFormer.
-To tackle these challenges, we introduce Kernel-U-Net, a symmetric and hierarchical U-shape neural network architecture.
-The kernel-U-Net encoder compresses gradually input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series.
+To tackle these challenges, we introduce Kernel-U-Net, a flexible and kernel-customizable U-shape neural network architecture.
+The kernel-U-Net encoder compresses the input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series.
 Specifically, Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels.
-Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear.
-Experiments on seven real-world datasets, considering both multivariate and univariate settings, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model PatchTST in the majority of cases and outperforms Yformer.
+Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; and 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear.
+Experiments on seven real-world datasets, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model in the majority of cases in channel-independent settings.
 The source code for Kernel-U-Net will be made publicly available for further research and application.
-1 Introduction
+publicationid: pubid: 979-8-3503-6813- 0/24/\$31.00 ©2024 IEEE
+1 Laboratoire Images, Signaux et Systèmes Intelligents (LISSI), Université Paris-Est Créteil , Île-de-France, France
+2 Département Informatique et Télécommunication, ESIEE Paris-Université Gustave Eiffel, Île-de-France, France
+3 HN-Services, Île-de-France, France
+4 Artificial Intelligence Laboratory, UMT, Tirana-Albanie
+Proceedings of the 18th International Conference On Innovations In Intelligent Systems And Applications, Craiova, Romania, 2024.
+I Introduction
 Time series forecasting predicts future trends based on recent historical information.
 It allows experts to track the incoming situation and react timely in critical cases.
-Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city Lai et al.
-(2018), monitoring influenza-like illness weekly patient cases Wu et al.
-(2021), monitoring electricity transformer temperature in the electric power long-term deployment Zhou et al.
-(2021) or forecasting temperature, pressure and humidity in weather station Liu et al.
-(2022b) etc.
-Over the past few decades, time series forecasting solutions have evolved from traditional statistical methodsKhandelwal et al.
-(2015) and machine learning techniquesPersson et al.
-(2017) to deep learning-based solutions, such as recurrent neural networks(RNN) Tokgöz and Ünal (2018), Long Short-term Memory (LSTM) Kong et al.
-(2017), Temporal Convolutional Network (TCN) Hewage et al.
-(2020) and Transformer-based model Li et al.
-(2019).
-Among the Transformer models applying to time series data, Informer Zhou et al.
-(2021), Autoformer Wu et al.
-(2021), and FEDformer Zhou et al.
-(2022) are the best variants that incrementally improved the quality of prediction.
-As a recent paper Zeng et al.
-(2023) challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in Nie et al.
-(2023) argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches.
+Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city [1], monitoring influenza-like illness weekly patient cases [2], monitoring electricity transformer temperature in the electric power long-term deployment [3] or forecasting temperature, pressure and humidity in weather station [4] etc.
+Figure 1: Illustration of Kernel U-Net Architecture.
+a) Architecture of Kernel U-Net, it allows executing linear kernel and nonlinear kernels such as MLP, LSTM, and Transformer.
+b) Illustration of K-U-Net Encoder, the application of $\phi^{(l)}_{enc}$ on patches is independent of the choice of kernel.
+c) In the K-U-Net decoder, a custom kernel $\phi^{(l)}_{dec}$ expands the vectors into patches in the reverse order.
+Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods[5] and machine learning techniques[6] to deep learning-based solutions, such as recurrent neural networks(RNN) [7], Long Short-term Memory (LSTM) [8], Temporal Convolutional Network (TCN) [9] and Transformer-based model [10].
+Among the Transformer models applying to time series data, Informer [3], Autoformer [2], and FEDformer [11] are the best variants that incrementally improved the quality of prediction.
+As a recent paper [12] challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in [13] argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches.
 By adding a linear patch layer, their model PatchTST successfully relieved the overfitting problem of transformer modules and reached state-of-the-art results.
@@
 Such integration should not only ensure a balance between computational efficiency and expressiveness but also respond to requests for rapid development and testing.
-The Convolutional U-net, as a classic and highly expressive model in medical image segmentationRonneberger et al.
-(2015), features a symmetric encoder and decoder structure that is elegant in its design.
+The Convolutional U-net, as a classic and highly expressive model in medical image segmentation[14], features a symmetric encoder and decoder structure that is elegant in its design.
 This model’s structure is particularly suited to the time series forecasting task, as both inputs and outputs in this context are typically derived from the same distribution.
-The first U-shape model adapted for time series forecasting was the YformerMadhusudhanan et al.
-(2023), which incorporated Transformers in both its encoder and decoder components.
+The first U-shape model adapted for time series forecasting was the Yformer[15], which incorporated Transformers in both its encoder and decoder components.
 As mentioned previously, employing Transformers on point-wise data has the potential to cause overfitting issues.
 Therefore, our investigation aims to discover if there is a U-shape architecture effective in time series forecasting, which also possesses the capability to integrate various modules flexibly, thus facilitating the specific customization of solutions.
-To tackle this problem, we propose a symmetric and hierarchical architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net and Yformer for time series forecasting.
-K-U-Net provides convenience for composing particular models with custom kernels that follow the design pattern.
-By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data.
+To tackle this problem, we propose a flexible and kernel-customizable architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net, Swin Transformer, and Yformer for time series forecasting.
+K-U-Net generalizes the concept of convolutional kernel and provides convenience for composing particular models with non-linear kernels.
+Following the design pattern, K-U-Net can easily integrate custom kernels by replacing linear kernels with Transformer or LSTM kernels.
+As a result, K-U-Net can gain expressivity by capturing more complex patterns and dependencies in the data.
 Furthermore, the hierarchical structure of K-U-Net exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning such sequences.
@@
 We compose 30 variants of K-U-Net by placing different kernels at different layers and then we choose the best candidates for fine-tuning.
-Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinearZeng et al.
-(2023) and PatchTST, in the majority of cases.
+Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinear[12] and PatchTST, in the majority of cases.
 In summary, the contributions of this work include:
-  • We propose Kernel-U-Net, a symmetric and hierarchical architecture that partitions input sequences into small patches at each layer of the network.
-  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the design pattern.
+  • We propose Kernel-U-Net, a U-shape architecture that progressively compresses the input sequence into a latent vector and expands it to generate the output sequence.
+  • Kernel-U-Net generalizes the concept of the convolutional kernel and provides convenience for composing particular models with custom kernels.
   • The computation complexity is guaranteed in linear when employing Transformer kernels at the second or higher layers.
@@
 We conclude that Kernel-U-Net stands out as a highly promising option for large-scale time series forecasting tasks.
-Its symmetric and hierarchical design provides a balance of low computational complexity and high expressiveness.
+Its hierarchical design provides a balance of low computational complexity and high expressiveness.
 In most scenarios, it either surpasses or is slightly below the state-of-the-art results.
 Furthermore, its adaptability in fast-paced development and testing environments is ensured by the use of flexible, customizable kernels.
-Figure 1: Comparison of the structures of different U-Net. a) classic Convolutional U-Net, b) Kernel U-Net, c) Swin Transformer U-Net
-2 Related works
-Transformer
-Transformer Vaswani et al.
-(2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks.
-It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.
+II Related works
+II-1 Transformer
+Transformer [16] was initially introduced in the field of Natural Language Processing (NLP) on language translation tasks.
+It contains a layer of positional encoding, blocks composed of layers of multiple head attentions, and a linear layer with SoftMax activation.
 As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route.
-Vision Transformers (ViTs) Dosovitskiy et al.
-(2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al.
-(2009).
-Swin Transformer Liu et al.
-(2021) proposed a hierarchical Transformer whose representation is computed with shifted windows.
+Vision Transformers (ViTs) [17] applied a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNet[18].
+Swin Transformer [19] proposed a hierarchical Transformer whose representation is computed with shifted windows.
 As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection.
@@
 In time series forecasting, the researchers were also attracted by transformer-based models.
-LogTrans Li et al.
-(2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer.
+LogTrans [10] proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer.
 To improve the computation efficiency, the authors propose also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck.
-Informer Zhou et al.
-(2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences.
+Informer [3] has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences.
 The authors propose a ProbSparse self-attention mechanism to replace the canonical self-attention efficiently.
 It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments.
-PyraformerLiu et al.
-(2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion.
+Pyraformer[20] simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion.
 Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass.
-PatchTST employed PropSparse and patches, thus reducing computation complexity to $O(\frac{L}{S}log(\frac{L}{S}))$, where $S$ is the size of the patch.
+PatchTST employs ProbSparse and a linear projection patch layer to reduce the computational complexity of the transformer to $O(\frac{L}{S}\log(\frac{L}{S}))$, where $S$ is patch size.
+However, its flattened head still incurs a computational cost of $O(\frac{L^{2}}{S})$.
 Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing.
-Autoformer Wu et al.
-(2021) introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods.
-Frequency Enhanced Decomposed Transformer (FEDformer)Zhou et al.
-(2022) captures global properties of time series with seasonal-trend decomposition.
+Autoformer [2] introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods.
+Frequency Enhanced Decomposed Transformer (FEDformer)[11] captures global properties of time series with seasonal-trend decomposition.
 The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
-U-Net
-U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al.
-(2015).
+II-2 U-Net
+U-Net is a neural network architecture designed primarily for medical image segmentation [14].
 U-net is composed of an encoder and a decoder.
-At the encoder phase, a long sequence is reduced by a convolutional layer and a max-pooling layer gradually into a latent vector.
+At the encoder phase, a long sequence is gradually reduced by a convolutional layer and a max-pooling layer into a latent vector.
 At the decoder stage, the latent vector is developed by a transposed convolutional layer for generating an output with the same shape as the input.
 With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
-With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al.
-(2015), biomedical 3D-Image segmentation Çiçek et al.
-(2016), time series segmentation Perslev et al.
-(2019), image super-resolution Han et al.
-(2022) and image denoising Zhang et al.
-(2023).
-Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al.
-(2016), 1D-U-Net Azar et al.
-(2022) and Swin-U-Net Cao et al.
-(2023) that only use transformer blocks at each layer of U-Net.
-In time series processing, U-time Perslev et al.
-(2019) is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer Madhusudhanan et al.
-(2023) is the first U-Net based model for time series forecasting task.
+With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation [14], biomedical 3D-Image segmentation [21], time series segmentation [22], image super-resolution [23] and image denoising [24].
+Its techniques evolve from basic 2D-U-Net to 3D U-Net [21], 1D-U-Net [25] and Swin-U-Net [26] that only use transformer blocks at each layer of U-Net.
+In time series processing, U-time [22] is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer [15] is the first U-Net based model for time series forecasting task.
 In particular, YFormer applied transformer blocks on each layer of U-Net and capitalized on multi-resolution feature maps effectively.
-Hierarchical and hybrid model
-In time series processing, the increasing size of data degrades the performance of deep models and increases crucially the cost of learning them.
+II-3 Hierarchical and hybrid model
+In time series processing, the increasing size of data degrades the performance of deep models and, crucially, increases the cost of learning them.
 For example, recurrent models such as RNN, LSTM, GRU have a linear complexity but suffer the gradient vanishing problem when input length increases.
@@
 To balance the expressiveness of complex models and computational efficiency, researchers investigated hybrid models that merge different modules into the network and hierarchical architectures.
-For example, authors in Du et al.
-(2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al.
-(2017) applied hierarchical LSTM and GRU for document classification.
-In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
-To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Duan et al.
-(2022), pyramidal structure in Pyraformer Liu et al.
-(2022a), AutoEncoders Asres et al.
-(2022), U-Net-like model in Yformer Madhusudhanan et al.
-(2023) or patches in PatchTST Nie et al.
-(2023).
-Kernel-U-Net is a symmetric and hierarchical U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences.
+For example, authors in [27] investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in [28] stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in [29] applied hierarchical LSTM and GRU for document classification.
+In time series processing, the authors in [30] combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
+To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Swin-Transformer [19], pyramidal structure in Pyraformer [20], U-shape structure in Yformer [15] or patch layer in PatchTST [13].
+Kernel-U-Net is a U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences.
 Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels.
@@
 Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
-3 Method
-3.1 Problem Formulation
-Let us note by $x\in R^{M\times N}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the size of the feature.
-Let L be the length of memory or the look-back window, so $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$ is a slice of length $L$ of all features.
-It contains historical information about the system at instant $t$.
-We call $X\in R^{N\times M\times L}$ the trajectory matrix which represents the sliced historical data for each instant $t\in[0,N-L-1]$.
-In the context of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$.
-Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window.
-Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps.
-Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
-3.2 Kernel U-Net
-Kernel U-Net(K-U-Net) is a hierarchical and symmetric U-shape architecture.
-It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels(Figure 1).
-More precisely, the K-U-Net encoder reshapes the input sequences into a large batch of small patches and repeatedly applies custom kernels on them until obtaining the latent vectors.
-Later, the K-U-Net decoder develops the batch of latent vectors into patches gradually at each layer and obtains a large batch of small patches(Figure 2).
-At last, K-U-Net reshapes the patches to get the final output.
-In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
-Figure 2: Hierarchical Partition
-Algorithm 1 Kernel Wrapper
-Input: kernel $f_{i}$, input patches $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
-Output: Instance of Kernel Wrapper
-  # Define the init function :
-  def __init(inputs)__:
-  # Initiation of main kernel function $f$
-  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
-  # Holds variables for global operation
-  _unet_skip_output = None
-  _unet_skip_input = None
-  # Define the forward function :
-  def forward(x):
-  reshape $x$ into (-1, $L_{in}$,$M_{in}$)
-  if _unet_skip_input is not None then
-           $x=x+\_unet\_skip\_input$
-  end if
-  # Apply $f$ on $x$
-  $y=f(x)$
-  _unet_skip_output = $y$
-  return $y$
-Hierarchical Partition of the input sequence.
-Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step.
-Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small patches $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all patches is equal to its original trajectory matrix:
-$X_{t}=Concat([[X_{t,i,j}\forall i\in[1,\frac{L}{L_{1}}]]\forall j\in[1,\frac{M}{M_{1}}]])$.
-The partitioned sequences form a large batch of patches and will be processed by kernels in the encoder.
-Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated patches $\hat{X}_{t+L,i,j}$.
-Hierarchical processing with kernels in the encoder and decoder
-Let us consider $X$ as a batch of trajectory matrix $X\in R^{B\times L\times M}$, where $B$ is the batch size.
-The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage.
-By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder.
-At the stage of the encoder, the shape of input $X$ is $(B,L,M)$.
-Let us note $L_{1}$ and $M_{1}$ as the length and feature size of a patch.
-Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem.
-We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of patches and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$.
-After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer.
-Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector (Algorithm 2).
-At the stage of the decoder, the operations are reversed.
-We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels.
-At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$.
-The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$ (Algorithm 3)..
-The algorithm is optimized in implementation because the for loop is expensive in the forward function of a PyTorch module.
-In our implementation, we nest the reshape and transpose operations in the kernel wrapper to avoid iteration over kernels.
-The complete algorithm is described in the next section.
-Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns.
-A kernel wrapper inherits ”nn.Module” class in Pytorch and offers an interface that executes the given kernel inside the encoder and decoder.
-It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector (Algorithm 1).
-By default, a kernel receives a batch of patches in the shape of $(-1,L_{i},M_{i})$ and outputs a new batch of patches in the shape of $(-1,L_{i+1},M_{i+1})$.
-Remark that the kernels reduce the length of patches from $L_{i}$ to $1$ in the encoder and expand the length of patches from $1$ to $L_{i}$ in the decoder.
-The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
-It reshapes the input patches and checks the output shape outside of the kernel.
-Algorithm 2 Kernel U-Net Encoder
-Input: $L_{1},M_{1}$, $[L_{2},$,...,… ,$L_{n},$ $L_{n+2},$,...,… ,$L_{n+m}]$, $[M_{2},$,...,… ,$M_{n},$ $M_{n+2},$,...,… ,$M_{n+m}]$, $[f_{1},f_{2},$,...,… ,$f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
+III Method
+III-A Problem Formulation
+Let us note by $x\in\mathbb{R}^{N\times M}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the feature size.
+Let $L$ be the length of memory or the look-back window, we denote the historical time series $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$.
+We also denote the future time series $(x_{t+L+1},...,x_{t+L+T})$, where $T$ is the length of future horizon and $t\in[0,N-L-T]$ is the time stamp.
+The time series forecasting task takes a multivariate time series as input and predicts a future series.
+Let $x_{t}$ be the features at the time step $t$, $L$ the length of the look-back window, and $T$ the future horizon.
+Given a historical data series $(x_{t+1},...,x_{t+L})$, time series forecasting task predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future.
+Then we can define the basic time series forecasting problem:
+$(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ (1)
+where $f$ is the function that predicts the future series based on a historical series.
+Algorithm A Kernel U-Net Encoder
+Input: $L_{1}$, $M_{1}$, $\{L_{2}$,...,$L_{n}$,$M_{2}$,...,$M_{m}\}$, $\{D_{2}$,...,$D_{n}$, $D_{n+2}$,..., $D_{n+m-1}\}$, $\{\phi^{(1)}$,...,$\phi^{(n)}$, $\phi^{(n+2)}$,...,$\phi^{(n+m)}\}$, $J_{h}$, $D_{h}$
 Output: Instance of Kernel U-Net Encoder
   # Define the init function :
-  def __init(inputs)__:
-  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
+  def __init__(inputs):
+  $J\in\{L_{1},L_{2},...,L_{n},$ $M_{2},...,M_{m}\}$
   # Create first layer
@@
 […]
-           layers.append(KernelWrapper($f_{i},$ $L_{i},$ $M_{i},$ $1,$
-           $M_{i+1}$))
+           layers.append(KW($\phi^{(l)},$ $J_{l},$ $D_{l},$ $1,$ $D_{l\_next}$))
   end for
   # Create last layer
-  layers.append(KernelWrapper($f_{n+m},L_{n+m},$
-  $M_{n+m},$ $L_{h},M_{h}$))
+  layers.append(KW($\phi^{(n+m)},J_{n+m},D_{n+m},J_{h},D_{h}$))
   encoder = nn.Sequential (*layers)
-Creation of Kernel U-Net
-Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of patch $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$ Remark that the hidden dimension of the model is the size of intermediate output vectors within layers.
-It corresponds to channel size in a convolutional network.
-This dimension can be augmented for a larger passage of information.
-We describe the creation of the K-U-Net encoder in Algorithm 2 and leave that of the decoder (Algorithm 3) and K-U-Net (Algorithm 4) in the appendix.
-3.3 Custom Kernels
-Linear kernel
-The linear kernel is a simple matrix multiplication (Figure 3). Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
-$Z_{i}=f_{i}(X)=X*w+b$
-, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector.
-Remark that the number of parameters of $w$ in a linear kernel is $L_{i}*1*ch_{in}*ch_{out}$ which is equivalent to that of a 1D convolutional layer whose kernel size is $L_{i}*1$.
-The multi-layer linear kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
-$Z_{i}=f_{i}(X)=Tanh(X*w_{1}+b_{1})*w_{2}+b_{2}$
-, where $Z_{i}$ is output of kernel, $w_{1}\in R^{L_{i}M_{i}\times L^{\prime}M^{\prime}}$ and $w_{1}\in R^{L^{\prime}M^{\prime}\times L_{i+1}M_{i+1}}$ are weight matrix , $b_{1}\in R^{L^{\prime}M^{\prime}}$ and $b_{2}\in R^{L_{i+1}M_{i+1}}$ are bias vectors, and $L^{\prime}=\frac{1}{2}(L_{i}+L_{i+1}),M^{\prime}=\frac{1}{2}(M_{i}+M_{i+1})$.
-Transformer kernel
-The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with Relu activation.
-The transformer kernel in this work follows the classical structure.
-LSTM kernel
-The LSTM kernel contains a classic LSTM cell and a linear layer for skip connection.
-The hidden states of all time steps are flattened and used for a linear combination in the next.
-The detailed structure of the Transformer kernel and the LSTM kernel is discussed in the appendix.
-Figure 3: Structure of Linear Kernel. a) Linear kernel, b) Multi-layer Linear Kernel with Tanh activation.
-Table 1: Multivariate time series forecasting results with Kernel U-Net.
+III-B Kernel U-Net
+Kernel U-Net (K-U-Net) is a neural network featuring hierarchical and symmetric U-shape architecture.
+It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels (Figure 1).
+More precisely, the K-U-Net encoder reshapes the input sequence into a large batch of small patches and repeatedly applies custom kernels on them until the latent vector is obtained.
+Later, the K-U-Net decoder expands the latent vector into patches gradually at each layer and obtains a large batch of small patches (Figure 1).
+At last, K-U-Net reshapes the patches to get the final output.
+In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
+III-B1 Hierarchical partition of the input and output
+In the first place, We split the input trajectory matrix into patches.
+Let us consider a trajectory matrix $X_{t}\in\mathbb{R}^{L\times M}$ at time step $t$.
+Given a list of multiples $\{L_{2},...,L_{n}\}$ for look-back window and $\{M_{2},...,M_{m}\}$ for feature, the patch length $L_{1}$ and feature unit $M_{1}$ such that $L=\prod_{1}^{n}L_{k}$ and $M=\prod_{1}^{m}M_{k}$.
+We reshape $X_{t}$ as a set of small patches $P_{t}=\{X_{t,i,j}|X_{t,i,j}\in\mathbb{R}^{L_{1}\times M_{1}}\}$, where $i\in\{1,...,\prod_{2}^{n}L_{k}\}$ and $j\in\{1,...,\prod_{2}^{m}M_{k}\}$.
+The total number of patches is the product of the multiples of length and feature size #$\#P_{t}=\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k}$.
+The partitioned patches will be processed by kernels in the encoder gradually and their size will be reduced after the kernel operation at each layer.
+In the decoder, the patches are generated from vectors of length $1$ at each layer.
+Since the decoder is symmetrical to the encoder, there will also be a trajectory matrix $\hat{X}_{t+L}\in\mathbb{R}^{T\times M}$ composed of a set of generated patches $\hat{P}_{t}=\{\hat{X}_{t+L,i,j}\}$ as output.
+For a simpler description, we let the look-back window $L$ and the forecasting horizon $T$ be equal.
+Algorithm B Kernel Wrapper (KW)
+Input: $\phi$, $X$, $J_{in}$, $D_{in}$, $J_{out}$, $D_{out}$
+Output: Instance of Kernel Wrapper
+  # Define the init function :
+  def __init__(inputs):
+  # Initiation of main kernel function $\phi$
+  ϕ=italic-$\phi=$ $\phi(J_{in},D_{in},J_{out},D_{out})$
+  # Holds variables for global operation
+  skip_out = None
+  skip_in = None
+  # Define the forward function :
+  def forward(x):
+  reshape $x$ into (-1, $J_{in}$,$D_{in}$)
+  $X=X+$ skip_in if skip_in is not None else $X$
+  $Z=\phi(X)$ assert $Z$.shape is (-1, $J_{out}$,$D_{out}$)
+  skip_out = $Z$ will be assigned to skip_in in decoder
+  return $Z$
+III-B2 Hierarchical processing with kernels
+The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage.
+By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder.
+Let us consider $\mathcal{X}\in\mathbb{R}^{B\times L\times M}$, a batch of trajectory matrix $X_{t}$, where $B$ is the batch size.
+We now describe the shape of the intermediate patch before and after the kernel operation.
+At the encoder stage, we compress input $\mathcal{X}\in\mathbb{R}^{B\times L\times M}$ into latent vector $\mathcal{Z}$.
+Let us assume that $D_{h}$ is the unique dimension of the hidden vectors at each intermediate layer and the latent vector to simplify the problem.
+Firstly, we reshape $\mathcal{X}$ to $(B,1,\prod_{2}^{n}L_{k},L_{1},\prod_{2}^{m}M_{k},M_{1})$ then transpose it to $(B,\prod_{2}^{m}M_{k},\prod_{2}^{n}L_{k},L_{1},1,M_{1})$ and reshape it to $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},L_{1},M_{1})$ We denote this vector $\mathcal{P}^{(1)}_{in}$ as a large batch of small patches ready for processing with a kernel.
+Secondly, the kernel at first layer can now process $\mathcal{P}^{(1)}_{in}$ and outputs a hidden vector $\mathcal{P}^{(1)}_{out}$ in shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},1,D_{h})$.
+After this operation, we reshape the output $\mathcal{P}^{(1)}_{out}$ to $\mathcal{P}^{(2)}_{in}$ of shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{3}^{n}L_{k},L_{2},D_{h})$ as input for the next layer.
+Iteratively, the encoder processes all the multiples $\{L_{2},...,L_{n},M_{2},...,M_{m}\}$ in order, and gives finally a batch of latent vector $\mathcal{Z}=\mathcal{P}^{(n+m)}_{out}$ in shape $(B,1,D_{h})$ (Algorithm A).
+TABLE I: Multivariate time series forecasting results with Kernel U-Net.
 The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
 We note the best results in bold and the second best results in underlined.
-Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer
+{tblr}
+cell11 = c=2, cell13 = c=2, cell15 = c=2, cell17 = c=2, cell19 = c=2, cell111 = c=2, cell113 = c=2, cell115 = c=2, cell117 = c=2, cell21 = c=2, cell31 = r=4, cell71 = r=4, cell111 = r=4, cell151 = r=4, cell191 = r=4, cell231 = r=4, cell271 = r=4, cell381 = c=2, cell383 = c=2, cell385 = c=2, cell387 = c=2, cell389 = c=2, cell3811 = c=2, cell3813 = c=2, cell3815 = c=2, cell391 = c=2, cell411 = r=4, cell451 = r=4, cell491 = r=4, cell531 = r=4, vline2-17 = 1, vline2-3,5,7,9,11,13,15,17 = -, hline1-3,7,11,15,19,23,27,31 = -, rowsep=0.12ex , colsep=3.25ex Methods & K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.370 0.393 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
-192 0.404 0.414 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
-336 0.420 0.430 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
-720 0.438 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
-ETTh2
-96 0.271 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
+ETTh1 96 0.355 0.388 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
+192 0.388 0.412 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
+336 0.407 0.427 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
+720 0.430 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
+ETTh2 96 0.269 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
 192 0.332 0.377 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931 1.593 1.021
-336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
-720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
-ETTm1
-96 0.286 0.342 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
-192 0.330 0.363 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
-336 0.360 0.384 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
-720 0.405 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
-ETTm2
-96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
-192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
-336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
-720 0.343 0.379 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
-Electricity
-96 0.129 0.226 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
-192 0.147 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
-336 0.163 0.261 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
-720 0.197 0.292 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
-Traffic
-96 0.358 0.253 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
-192 0.373 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
-336 0.390 0.271 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
-720 0.430 0.292 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
-Weather
-96 0.142 0.195 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
-192 0.188 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
-336 0.241 0.284 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
-720 0.310 0.333 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
-3.4 Complexity Analysis
-As the K-U-Net is symmetric, we study the complexity of the encoder.
+336 0.355 0.400 0.329 0.384 0.357 0.400 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
+720 0.384 0.435 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
+ETTm1 96 0.275 0.331 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
+192 0.320 0.361 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
+336 0.349 0.380 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
+720 0.401 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
+ETTm2 96 0.157 0.243 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
+192 0.213 0.283 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
+336 0.266 0.320 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
+720 0.343 0.377 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
+Electricity 96 0.128 0.219 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
+192 0.145 0.234 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
+336 0.160 0.250 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
+720 0.196 0.283 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
+Traffic 96 0.354 0.229 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
+192 0.372 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
+336 0.388 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
+720 0.430 0.269 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
+Weather 96 0.142 0.183 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
+192 0.187 0.226 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
+336 0.238 0.269 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
+720 0.308 0.323 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
+At the stage of the decoder, the operations are reversed.
+We start with $\mathcal{Q}^{(n+m)}_{in}=\mathcal{Z}$ of shape $(B,1,D_{h})$, a set of input patches to the decoder.
+We send it to decoder kernel and get an output $\mathcal{Q}^{(n+m)}_{out}$ in shape $(B,M_{m},D_{h})$, then we reshape it to $\mathcal{Q}^{(n+m-1)}_{in}$ in shape $(B\cdot M_{m},1,D_{h})$ for the next kernel.
+At the end of this iteration over multiples $,\{M_{m-1},$...,… , $M_{2},$ $L_{n},$ ,...,… , $L_{2},$ $L_{1}\}$, we finally have a set of patches $\mathcal{Q}^{(1)}_{out}$ in the shape $(B\cdot\prod_{2}^{m}M_{k}\cdot\prod_{2}^{n}L_{k},L_{1},M_{1})$.
+The last operations are reshaping it into $(B,\prod_{2}^{m}M_{k},\prod_{2}^{n}L_{k},L_{1},1,M_{1})$ , transposing it into $(B,1,\prod_{2}^{n}L_{k},L_{1},\prod_{2}^{m}M_{k},M_{1})$ and reshaping it into $(B,L,M)$ as final output.
+III-B3 Kernel Wrapper
+The Kernel Wrapper (KW) requires necessary parameters such as kernel $\phi$, input patches $X$, output patches $Z$, patches set size $\mathcal{B}$, input patch length $J_{in}$ and dimension $D_{in}$, output patch length $J_{out}$ and dimension $D_{out}$.
+The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
+It reshapes the input patches, executes with the kernel inside, and then checks the output shape.
+The wrapper processes $X$ and outputs $Z$ in the encoder, or it processes the sum of $X$ and an encoder output via skip connection, then gives an output in the decoder (Algorithm B).
+III-B4 Formulation of Kernel Operation
+We add enc and dec in the index of variables to differentiate their utilization in the encoder and decoder.
+We denote $\mathcal{P}=X_{enc}$ and $\mathcal{Q}=X_{dec}$ the set of patches, $l\in\{1,...,n,n+2,...,n+m\}$ the index of layers, and $\mathcal{B}$ the patches set size.
+By default, an encoder kernel at layer $l$ receives $\mathcal{P}^{(l)}_{in}\in\mathbb{R}^{\mathcal{B}^{(l)}\times J^{(l)}_{enc\_in}\times D^{(l)}_{enc\_in}}$ and gives $\mathcal{P}^{(l)}_{out}\in\mathbb{R}^{\mathcal{B}^{(l)}\times 1\times D^{(l)}_{enc\_out}}$.
+The decoder kernel at layer $l$ receives $\mathcal{Q}^{(l)}_{in}\in\mathbb{R}^{\mathcal{B}^{(l)}\times 1\times D^{(l)}_{dec\_in}}$ and outputs $\mathcal{Q}^{(l)}_{out}\in\mathbb{R}^{\mathcal{B}^{(l)}\times J^{(l)}_{dec\_out}\times D^{(l)}_{dec\_out}}$.
+We recall that $i^{(l)},j^{(l)}$ are the indices for multiples of length and feature at layer $l$.
+Given a set of input $\mathcal{P}^{(l)}_{in}$ at layer $l$, the output $\mathcal{P}^{(l)}_{out}$ of kernel $\phi^{(l)}_{enc}$ in encoder is written :
+$\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}=\phi^{(l)}_{enc}(\mathcal{P}^{(l)}_{in,i^{(l)},j^{(l)}})$ (2)
+The decoder kernel $\phi^{(l)}_{dec}$ at layer $l$ takes the sum of input $\mathcal{Q}^{(l)}_{in}$ and encoder output $\mathcal{P}^{(l)}_{out}$ via the skip connection as input and produces $\mathcal{Q}^{(l)}_{out}$ as output:
+$\mathcal{Q}^{(l)}_{out,i^{(l)},j^{(l)}}=\phi^{(l)}_{dec}(\mathcal{Q}^{(l)}_{in,i^{(l)},j^{(l)}}+\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}})$ (3)
+Remark that in case of $l=n+m$, we have $\mathcal{Q}^{(l)}_{in,i^{(l)},j^{(l)}}+\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}=\mathcal{P}^{(l)}_{out,i^{(l)},j^{(l)}}$ because there are no higher layers and the kernel only process the encoder output.
+III-B5 Creation of Kernel U-Net
+We create the encoder, decoder, and the U-Net in order.
+The algorithm passes parameters describing the multiples, kernels, and hidden dimensions.
+Let us note the input length $L$, feature dimension $M$, concatenated lists of multiples of look-back window and feature $,\{L_{2},$...,… , $L_{n},$ $M_{2},$ ,...,… , $M_{m}\}$, list of hidden dimension of same size $,\{D_{h},$...,… , $D_{h}\}$, patch size $L_{1}$ and feature unit $M_{1}$ , a list of kernels $\{\phi^{(1)}_{enc},$ $\phi^{(2)}_{enc},$ ,...,… , $\phi^{(n)}_{enc},$ $\phi^{(n+2)}_{enc},$ ,...,… , $\phi^{(n+m)}_{enc}\}$, latent vector length $J_{h}$ and dimension $D_{h}$.
+We also use next and prev to iterate over the index $l$.
+We set the hidden dimension of intermediate output vectors within layers to be equal to that of latent vector for simpler description.
+It corresponds to channel size in a convolutional network and can be augmented for a larger passage of information if necessary.
+We describe the creation of the K-U-Net encoder in Algorithm A.
+The decoder is symmetrical to the encoder and applies kernels in reverse order.
+More precisely, the decoder takes $J_{h}$ and $D_{h}$, multiples $J\in\{L_{1},$ $L_{2},...,$ $L_{n},$ $M_{2},$ ,...,… , $M_{m}\}$, kernels $,\{\phi^{(n+m)}_{dec},$...,… , $\phi^{(n+2)}_{dec},$ $\phi^{(n)}_{dec},$ ,...,… , $\phi^{(2)}_{dec},$ $\phi^{(1)}_{dec}\}$ and initiates kernel wrappers KW($\phi^{(n+m)}_{dec},$ $J_{h},$ $D_{h},$ $J_{n+m},D_{n+m})$, KW($\phi^{(l)}_{dec},$ $1,$ $D_{l\_prev},$ $J_{l},D_{l})$, KW($\phi^{(1)}_{dec},$ $1,$ $D_{2},$ $L_{1},M_{1})$ at the highest, intermediate, and lowest layers respectively.
+The K-U-Net initiates an encoder and a decoder.
+In the forward function, the encoder processes the input and generates a list of outputs at each layer and a latent vector.
+It assigns the skip_out from encoder kernels to skip_in in decoder kernels.
+Then the decoder takes the outputs from the encoder via skip-connection and the latent vector to generate the final result.
+Figure 2: Structure of Linear Kernel. a) Linear Kernel, b) Multi-layer perceptron (MLP) Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
+III-C Custom Kernels
+III-C1 Linear kernel
+The linear kernel is a simple matrix multiplication.
+Given $X\in\mathbb{R}^{\mathcal{B}\times J_{in}\times D_{in}}$, linear kernel $\phi$ reshape it to $(\mathcal{B},J_{in}\cdot D_{in})$ and process it as follow:
+$Z=\phi(X)=XW+b$ (4)
+where $Z$ is output of kernel, $W\in\mathbb{R}^{J_{in}\cdot D_{in}\times J_{out}\cdot D_{out}}$ is weight matrix and $b\in\mathbb{R}^{J_{out}\cdot D_{out}}$ is bias vector.
+Remark that the number of parameters of $W$ is $J_{in}\cdot D_{in}\cdot J_{out}\cdot D_{out}$ and this kernel operation is equivalent to the process with a 1D convolutional layer whose kernel size is $J_{in}$ or $J_{out}$.
+III-C2 Multi-Layer Perceptron kernel
+The multi-layer perceptron (MLP) kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
+$Z=\phi(X)=Tanh(XW_{1}+b_{1})W_{2}+b_{2}$ (5)
+where $Z$ is output of kernel, $W_{1}\in\mathbb{R}^{J_{in}\cdot D_{in}\times J^{\prime}\cdot D^{\prime}}$ and $W_{2}\in\mathbb{R}^{J^{\prime}\cdot D^{\prime}\times J_{out}\cdot D_{out}}$ are weight matrices , $b_{1}\in\mathbb{R}^{J^{\prime}\cdot D^{\prime}}$ and $b_{2}\in\mathbb{R}^{J_{out}\cdot D_{out}}$ are bias vectors, and $J^{\prime}=\frac{1}{2}(J_{in}+J_{out}),D^{\prime}=\frac{1}{2}(D_{in}+D_{out})$.
+III-C3 Transformer kernel
+The Vanilla Transformer[16] is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with ReLU activation.
+The transformer kernel in this work follows the classical structure (Figure 2).
+III-C4 LSTM kernel
+The LSTM kernel contains a classic LSTM cell[31] and a linear layer for skip connection. The hidden states of all time steps are combined with a linear layer in the next (Figure 2).
+III-D Complexity Analysis
+Since the K-U-Net is symmetric, we only study the complexity of the encoder layer and assume that the feature size is $1$.
 Let us suppose that a kernel is receiving a sequence of length $L$ where $L=\prod_{1}^{n}L_{i}$ and all $L_{i}$ are equal.
 As the patch size in the first layer of Kernel U-Net encoder is $L_{1}$, a kernel module will process $\frac{L}{L_{1}}$ patches of size $L_{1}$.
-Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}*f(L_{1}))$ where $f(L_{1})$ is the complexity inside the kernel in the function of the patch size $L_{1}$ and $n$ is the index of the layer.
-In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}*L_{1})$= $O(L)$.
-In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}*L_{1}^{2})$= $O(L*L_{1})$.
-Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(L*log(L))$.
-Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}^{2}}*L_{1}^{2})$= $O(L)$.
-Following the same demonstration, the complexity of using multi-layer linear and LSTM kernels starting from the second layer is also bounded by $O(L)$.
-Figure 4: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on ETTh1 dataset.
-4 Experiments and Results
-Datasets.
-We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^1] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Traffic [^3] http://pems.dot.ca.gov and Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams
-20112014
-. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
-Figure 5: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on Weather dataset.
-Here, we followed the experiment setting in Zeng et al.
-(2023) and partitioned the data into $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset.
-The data is split into $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
-Baselines and Experimental Settings.
-We follow the experiment setting in NLinear Zeng et al.
-(2023) and take $L=\{336,720\}$ step historical data as input then forecast $T\in\{96,192,336,720\}$ step value in the future.
-We replace the last value normalization with mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalizationUlyanov et al.
-(2017) for the Traffic dataset.
-We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as mentioned in Wu et al.
-(2021).
-We include recent methods: PatchTSTNie et al.
-(2023), NLinear, DLinearZeng et al.
-(2023), FEDformer Zhou et al.
-(2022), Autoformer Wu et al.
-(2021), Informer Zhou et al.
-(2021), LogTrans Li et al.
-(2019) YformerMadhusudhanan et al.
-(2023).
-We merge the result reported in Nie et al.
-(2023) and Zeng et al.
-(2023) for taking their best in a supervised setting and execute Yformer in our environment with default parameters in its Github [^5] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
-Figure 6: Computation time of K-U-Net, PatchTST and Yformer.
-Model Variants Search.
+Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}\cdot g(L_{1}))$ where $g(L_{1})$ is the complexity inside the kernel in the function of patch size $L_{1}$ and $n$ is the index of the layer.
+In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}\cdot L_{1})$= $O(L)$.
+In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}\cdot L_{1}^{2})$= $O(L\cdot L_{1})$.
+Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(Llog(L))$.
+Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}\cdot L_{2}}\cdot L_{2}^{2})$= $O(L)$.
+Following the same demonstration, the complexity of using LSTM kernels and MLP starting from the second layer is also bounded by $O(L)$.
+IV Experiments and Results
+IV-1 Datasets
+We conducted experiments with our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^2] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^3] https://www.bgc-jena.mpg.de/wetter/, Traffic [^4] http://pems.dot.ca.gov and Electricity [^5] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014.
+These datasets have been benchmarked and publicly available on[2] and the description of the dataset is available in [12].
+Here, we followed the experiment setting in [12] and partitioned the data to $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset.
+The data is split to $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
+IV-2 Baselines and Experimental Settings
+We follow the experiment setting in NLinear [12] and take $L$ step historical data as input then forecast $T$ step value in the future where $L\in\mathcal{L}=\{336,720\}$ and $T\in\mathcal{T}=\{96,192,336,720\}$.
+We replace the last value normalization by mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalization[32] for the Traffic dataset.
+We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) for evaluation as mentioned in [2].
+We include recent methods: PatchTST[13], NLinear, DLinear[12], FEDformer [11], Autoformer [2], Informer [3], LogTrans [10] Yformer[15].
+We merge the result reported in [13] and [12] for taking their best in a supervised setting and execute Yformer with default parameters in its Github [^6] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
+IV-3 Experiment details
+We use a 4-layer K-U-Net for experiments.
+The list of multiples are respectively {4, 3, 7} and {6, 6, 5} for look-back windows $L\in\{336,720\}$.
+The bottom patch length is 4 and its width is 1.
+We reshape the input into $(B\cdot M,L,1)$ for processing with K-U-Net as we follow the channel-independent setting [12] and then reshape it back.
+The hidden dimensions are 128 for multivariate tasks.
+The learning rate is selected in $[0.00005,0.001]$.
+The training epoch is 50 and the patience of early stopping is 10 in general.
+We use MAE as the loss function.
+IV-4 Model Variants Search
 We propose 4 kernels for experiments with K-U-Net on 7 datasets.
-By replacing a linear kernel at different layers with other types of kernels for searching for variants that adapt the dataset.
-The variants are noted as ”K-U-Net $<$kernel$>$_$<$replacing_layer$>$ ($<$lookback_window$>$)”.
-For example, K-U-Net Linear_$\_0000(720)$ means that the model is made of linear kernels at all layers and the lookback window size is 720, K-U-Net Transf_$\_0100(720)$ means that a transformer kernel replaces the linear kernel at the second layer, K-U-Net Linear_$\_0110(720)$ means that multilayer linear kernels replace the kernels at the second and third layers.
-We compose 16 variants with Multi-layer linear kernels and 7 variants with Transformer and LSTM kernels.
+By replacing a linear kernel at different layers with other types of kernels, we search for variants that adapt the dataset.
+As there are too many variants to name, we note them as ”KUN_$<$kernel$>$_$<$replace_layer$>$ ($<$look-back_window$>$)”.
+For example, KUN_Linear_$\_0000$ $(720)$ means that the model is made of linear kernels at all layers and the look-back window size is 720, KUN_Transf_$\_0100$ means that a transformer kernel replaces the linear kernel at the second layer, KUN_Linear_$\_0110$ means that multilayer perceptron kernels replace the kernels at the second and third layers.
+We have composed 16 variants with MLP kernels and 7 variants with Transformer and LSTM kernels.
+Figure 3: Search result of K-U-Net with Linear, MLP, LSTM, Transformer kernels on ETTh1 dataset.
+Figure 4: GPU consumation of K-U-Net, PatchTST and Yformer.
 To enumerate all variants that achieve at least once the best result, we report their performance by the average of the top 5 minimum running MSE (Top5MMSE) values on the validation and test set.
 The search results are reported with a relative score(RS) to the minimum Top5-MMSE:
-$RS(m)=Min_{t\in T}\frac{Top5MMSE(m,t)}{Min_{\{m\in M,l\in L\}}Top5MMSE(m,t,l)}$
-, where $m$ is a model in the models set $M$, $t$ is a forecasting horizon in $T$, $l$ is a look back window in $L$.
+$\text{RS}(V)=\text{Min}_{T\in\mathcal{T}}\frac{\text{Top5MMSE}(V,T)}{\text{Min}_{\{U\in\mathcal{V},L\in\mathcal{L}\}}\text{Top5MMSE}(U,T,L)}$ (6)
+, where $U$, $V$ are variant models in the models set $\mathcal{V}$, $T$ is a forecasting horizon in $\mathcal{T}$, $L$ is a look back window in $\mathcal{L}$.
 Relative score notes the best-performed model with 1 and thus helps to identify the high-potential candidates for further fine-tuning examinations.
-We observe in Figure 4 that the best variant for ETTh1 dataset is K-U-Net Linear_$\_0010$.
-Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting.
-Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are K-U-Net Linear_$\_0110$, K-U-Net LSTM_$\_0100$ and K-U-Net Transf_$\_0100$.
-Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
+We observe in Figure 3 that the best variant for ETTh1 dataset is KUN_Linear_$\_0010$.
+Comparing the relative score of the K-U-Net of replace__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting.
+Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are KUN_Linear_$\_0110$, KUN_LSTM_$\_0100$ and KUN_Transf_$\_0100$.
+Comparing the relative score of the K-U-Net of replace__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
 Among candidates in the search phase, we empirically choose 3 variants for fine-tuning experiments with 5 runs.
-The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in univariate settings and multivariate settings in most cases.
-Multivariate time series forecasting result.
-We remark that our model improves the MSE performance around $72.92\%$ compared with Yformer and $0.99\%$ compared to PatchTST and NLinear in the multivariate setting (Table 1).
-It is worth noting that K-U-Net achieves similar results on the Electricity dataset with a variant based on Multilayer linear kernels.
-Univariate time series forecasting result.
-Additionally, we remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3).
-It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ compared to PatchTST.
-Figure 7: GPU consumation of K-U-Net, PatchTST and Yformer.
-Computation Efficiency
-We examined the computation efficiency of 4 variants of K-U-Net, PatchTST, and Yformer.
+The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in multivariate settings in most cases.
+IV-5 Multivariate time series forecasting result
+We remark that our model improves the MSE performance around $72.92\%$ compared with Yformer and $2.99\%$ compared to PatchTST and NLinear in the multivariate setting (Table I).
+It is worth noting that K-U-Net achieves similar results on the Electricity dataset with a variant based on MLP kernels.
+Figure 5: Search result of K-U-Net with linear, MLP, LSTM, Transformer kernels on Weather dataset.
+Figure 6: Computation time of K-U-Net, PatchTST and Yformer.
+IV-6 Computation Efficiency
+We examined the computation efficiency of 6 variants of K-U-Net, PatchTST, and Yformer.
 We execute the models on the ETTh1 and ETTm1 datasets for 10 epochs and measure the average execution time per epoch and the GPU consumption during the training.
 The hidden dimension is set to 128 for all models.
-For fair comparison, PatchTST and Yformer contain 2 layers of Transformer block which equals to K-U-Net_Transf_$\_0100$.
+For fair comparison, PatchTST and Yformer contain 2 layers of Transformer block which equals to KUN_Transf_$\_0100$.
 All experiments are executed on a Tesla V100 GPU in a Google Colab environment.
-Comparing to the PatchTST, K-U-Net Linear_$\_0000$ saves $85.20\%$ and $83.07\%$ memory (Figure 7) and saves $81.82\%$ and $80.55\%$ computation time (Figure 6) on ETTh1 and ETTm1 datasets, K-U-Net_Transf_$\_0100$ saves $55.15\%$ and $50.00\%$ memory and saves $28.47\%$ and $26.88\%$ computation time respectively.
-5 Conclusion
+Comparing to the PatchTST, KUN_Linear_$\_0000$ saves $85.20\%$ and $83.07\%$ memory (Figure 4) and saves $81.82\%$ and $80.55\%$ computation time (Figure 6) on ETTh1 and ETTm1 datasets, KUN_Transf_$\_0100$ saves $55.15\%$ and $50.00\%$ memory and saves $28.47\%$ and $26.88\%$ computation time respectively.
+V Conclusion
 In this paper, we propose Kernel-U-Net, a highly potential candidate for large-scale time series forecasting tasks.
@@
 In the future, we hope to develop more kernels and hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.
-Appendix A Appendix
-A.1 Complement Algorithms
-The complete algorithm of Kernel U-Net contains creation of Encoder, Decoder (Algorithm 3), Wrapper, and Kernel U-Net(Algorithm 4).
-Algorithm 3 Kernel U-Net Decoder
-Input: $L_{h}$, $M_{h}$, $[L_{n},...,L_{2}]$, $[L_{n+m},...,L_{n+2}]$, $L_{1}$, $M_{1}$, $[f_{n+m},...,f_{n+2},$ $f_{n},...,f_{2}]$, $[M_{n+m},...,M_{n+2},$ $M_{n},...,M_{2}]$, $L_{1}$, $M_{1}$.
-Output: Instance of Kernel U-Net Decoder
-  # Define the init function :
-  def __init(inputs)__:
-  # Create a list of multiples
-[…]
-  # Create first layer
-  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$
-  $L_{n+m},$ $M_{n+m}$)]
-  # Create intermediate layers
-  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
-           layers.append(KernelWrapper($f_{i},$ $1,M_{i+1}$
-           $,L_{i},M_{i})$)
-  end for
-  # Create last layer
-  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
-  decoder = nn.Sequential (*layers)
-Algorithm 4 Kernel U-Net
-Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
-Output: Instance of Kernel U-Net
-  # Define the init function :
-  def __init(inputs)__:
-  # Initiation of encoder and decoder in K-U-N.
-  encoder = KernelUNetEncoder(inputs)
-  decoder = KernelUNetDecoder(inputs)
-  # Define the forward function :
-  def forward(x):
-  reshape x to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
-  transpose x to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
-  reshape x to $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$
-  # Apply encoder on x
-  z = encoder(x)
-  for  $i$ in $[1,...,n]$ do
-           output = encoder.layers[i]._unet_skip_output
-           decoder.layers[$n-i+1$]._unet_skip_input
-           = output
-  end for
-  # Apply decoder on z
-  y= decoder(z)
-  reshape y to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
-  transpose y to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
-  reshape y to $(B,L,M)$
-  return y
-A.2 LSTM and Transformer kernels
-Transformer kernel
-The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.
-We follow the description in Wen et al.
-[2023].
-Positional encoding (PE).
-In a vanilla Transformer, the input sequence is added by a positional encoding vector given by:
-$PE(t)=\left\{\begin{array}[]{ll}sin(\omega it)&ifi\%2=0\\ cos(\omega it)&ifi\%2=1\\ \end{array}\right.$
-, where $\omega i$ is the frequency of waves generated at each dimension.
-By adding positional information to the input, the transformer can consider the same patterns that occur in different time steps as different patterns.
-Figure 8: Structure of Kernels a) Linear Kernel, b) Multi layer linear Kernel with Tanh activation, c) LSTM Kernel, d) Transformer Kernel
-Multiple head attentions.
-With the Query-Key-Value $(QKV)$ model, the scaled dot-product attention used by Transformer is given by
-$Attention(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{D_{k}}}))V$
-, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values.
-The multiple attention block contains repeated heads and can be given by
-$MultiHeadAttn(Q,K,V)=Concat(h_{1},...,h_{H})W^{O}$
-, where $h_{i}=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}).$
-The transformer kernel in this work (Figure 8) follows the classical structure with 2 additional linear layers.
-The first linear layer before the Transformer block of size $M_{in}\times M_{out}$ reduces the dimension.
-The last linear layer is a matrix of size $L_{in}*M_{out}\times L_{out}*M_{out}$ that reduces the input length.
-LSTM kernel
-The Long Short-term Memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$.
-Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$.
-The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$.
-The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state.
-The formulation of the LSTM cell is as follows:
-$f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$
-$i_{t}=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$
-$C^{\prime}_{t}=tanh(W_{C}[h_{t-1},x_{t}]+b_{C})$
-$C_{t}=f_{t}C_{t-1}+i_{t}C^{\prime}_{t}$
-$o_{t}=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$
-$h_{t}=o_{t}tanh(C_{t})$
-In our experiments, we add linear layers (Figure 8) that take the flattened output of all hidden states of the LSTM cell for linear combination.
-$Z=W*Flatten([h_{1},...,h_{t}])+b$
-, where $W\in R^{L_{i}*d_{h}}$, $b\in R^{L_{i+1}*M_{i+1}}$ and $d_{h}$ is the dimension of hidden state and $d_{h}=M_{i}$ in general.
-The residual skip operation aside from the LSTM cell contains a linear layer equivalent to a simple linear kernel discussed in the previous sections.
-A.3 Dataset and Experiment
-Dataset Statistics
-We use 7 popular multivariate datasets for experiments.
-We create subset Weather(S), Electricity(S), and Traffic(S) from top-left to bottom-right for quick search of K-U-Net variants.
-The statistic is in noted in Table 2.
-We refer the description of datasets in Nie et al.
-[2023]:
-  • Weather dataset collects 21 meteorological indicators in Germany, such as humidity and air temperature.
-  • Traffic dataset records the road occupancy rates from different sensors on San Francisco freeways.
-  • Electricity is a dataset that describes 321 customers’ hourly electricity consumption.
-  • ETT (Electricity Transformer Temperature) datasets are collected from two different electric transformers labeled with 1 and 2, and each of them contains 2 different resolutions (15 minutes and 1 hour) denoted with m and h.
-Table 2: Statistics of popular datasets for experiments.
-Datasets ETTh1 ETTh2
-ETTm1
-ETTm2 Weather
-Traffic
-Electricity Weather(S) Traffic(S) Electricity(S)
-Features 7 7
-7
-7 21
-862
-321 21 (100%) 21 (2.5%) 22 (7%)
-Timesteps 17420 17420
-69680
-69680 52696
-17544
-26304 10539 (20%) 12280 (70%) 10521 (40%)
-Table 3: Univariate time series forecasting results with Kernel U-Net.
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold and the second best in underlined.
-Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer LogTrans Yformer
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.053 0.177 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468 0.211 0.379
-192 0.069 0.205 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409 0.228 0.403
-336 0.076 0.221 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546 0.179 0.355
-720 0.077 0.224 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629 0.260 0.444
-ETTh2
-96 0.114 0.267 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379 0.240 0.398
-192 0.146 0.306 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429 0.270 0.429
-336 0.157 0.321 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437 0.263 0.422
-720 0.182 0.347 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387 0.264 0.423
-ETTm1
-96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171 0.211 0.393
-192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317 0.262 0.42
-336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459 0.424 0.559
-720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579 0.413 0.552
-ETTm2
-96 0.062 0.182 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208 0.128 0.275
-192 0.089 0.224 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275 0.155 0.308
-336 0.114 0.259 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302 0.208 0.358
-720 0.156 0.312 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321 0.261 0.405
-Table 4: Comparison of 16 variants of Linear K-U-Net in baseline setting and without using Skip connection(-Skip), Random Erase(-RE), Chanel Independence(-CI).
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold and the second best results in underlined for minimum score(Min) and average score(Mean).
-Method Min (Base) Min (-Skip) Min (-RE) Min (-CI) Mean (Base) Mean (-Skip) Mean (-RE) Mean (-CI)
-Metric MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE MSE Top5-
-MMSE
-ETTh1
-96 0.368 0.368 0.375 0.375 0.370 0.369 0.459 0.451 0.376 0.374 0.393 0.385 0.377 0.373 0.488 0.483
-192 0.404 0.402 0.412 0.413 0.402 0.402 0.506 0.495 0.409 0.407 0.423 0.423 0.411 0.407 0.551 0.538
-336 0.423 0.421 0.443 0.441 0.427 0.422 0.535 0.519 0.443 0.430 0.489 0.455 0.442 0.431 0.581 0.549
-720 0.457 0.439 0.461 0.453 0.461 0.438 0.629 0.576 0.492 0.472 0.531 0.515 0.488 0.470 0.690 0.598
-ETTh2
-96 0.274 0.272 0.273 0.273 0.274 0.273 0.423 0.397 0.302 0.287 0.299 0.290 0.300 0.286 0.458 0.415
-192 0.336 0.353 0.337 0.355 0.336 0.354 0.414 0.442 0.370 0.375 0.376 0.382 0.372 0.374 0.445 0.474
-336 0.361 0.361 0.364 0.364 0.362 0.361 0.412 0.422 0.402 0.384 0.404 0.392 0.399 0.384 0.428 0.444
-720 0.398 0.394 0.397 0.397 0.395 0.394 0.453 0.456 0.451 0.418 0.459 0.436 0.444 0.420 0.468 0.519
-ETTm1
-96 0.288 0.284 0.289 0.290 0.288 0.284 0.321 0.318 0.296 0.292 0.303 0.298 0.296 0.291 0.346 0.333
-192 0.324 0.327 0.325 0.324 0.326 0.324 0.359 0.358 0.335 0.331 0.341 0.336 0.334 0.331 0.383 0.373
-336 0.359 0.358 0.353 0.354 0.359 0.357 0.391 0.391 0.366 0.363 0.374 0.368 0.366 0.362 0.418 0.406
-720 0.408 0.404 0.403 0.403 0.409 0.403 0.452 0.447 0.418 0.411 0.419 0.413 0.418 0.410 0.478 0.467
-ETTm2
-96 0.162 0.162 0.162 0.161 0.163 0.161 0.213 0.204 0.168 0.165 0.170 0.165 0.168 0.164 0.228 0.210
-192 0.218 0.217 0.216 0.215 0.217 0.216 0.298 0.274 0.228 0.223 0.229 0.224 0.228 0.222 0.318 0.286
-336 0.270 0.268 0.268 0.267 0.265 0.265 0.372 0.345 0.283 0.277 0.281 0.277 0.284 0.272 0.394 0.356
-720 0.342 0.349 0.343 0.347 0.343 0.347 0.410 0.408 0.354 0.361 0.351 0.361 0.364 0.356 0.434 0.421
-Average decrease +0.42% +0.77% +0.03% -0.20% +24.19% +21.45% +2.24% +2.30% +0.03% -0.32% +25.92% +23.72%
-Experiment details
-We use a 4-layer K-U-Net for experiments.
-The list of multiples are respectively [4, 4, 3, 7] and [4, 6, 6, 5] for look-back windows $L$ and forecasting horizons $T$ in {336, 720}.
-Remark that $T$ is fixed to be 336 for shorter forecasting cases.
-The bottom patch length is 4 and its width is 1 as we follow the Channel-Independent (CI) setting in Zeng et al.
-[2023].
-In case of not using CI, the patch width equals to feature size.
-The hidden dimensions are 128 for multivariate and univariate tasks in general, except that they are 256 on the Traffic dataset.
-The learning rate is selected in $[0.00005,0.001]$ depending on the dataset.
-The training epoch is 50 and the patience of early stopping is 10 in general.
-We apply weighted early stopping where the triggering MSE (TMSE) is computed with $TMSE_{e+1}=\tau*TMSE_{e}+(1-\tau)*VMSE_{e}$ , where $\tau$ is a hyper-parameter to be set, $e$ is the index of the episode, and VMSE is the running MSE on a validation set.
-The parameter $\tau$ sets the sensitivity of early stopping and is empirically selected.
-In general, it is set to be 0.9 in the variants searching stage and 0.5 in the fine-tuning stage.
-The MSE loss function is defined with $:\mathcal{L}=\mathbb{E}_{\boldsymbol{x}}\frac{1}{B*M}\sum_{i=1}^{B*M}\left\|\hat{\boldsymbol{x}}_{L+1:L+T}^{(i)}-\boldsymbol{x}_{L+1:L+T}^{(i)}\right\|_{2}^{2}$ , where $B$ is the batch size and $M$ is the multiple of feature size unit of time series.
-A.4 Search of model variants
-We searched 30 variants of K-U-Net across Linear_[$0$,…,$1$] and {LSTM, Transformer}_[0100,…,$1$] for the ETT, Weather datasets and 3 subsets.
-The lookback window is typically 720, and we additionally tested a version of 336 on the ETT dataset.
-The result is reported with the average of running Top5 Minimum MSE (Top5MMSE) on the validation and test set in a single run.
-This measures the extreme performance of a variant and its potential in the fine-tuning stage.
-We use relative score (RS) to locate best-performed variants (i.g.
-$RS=1$ means that it’s the best among all candidates).
-We plot the RS of variants of 5 datasets in Figures 9 and Figures 10.
-Based on the search result, We note that Linear and Multilayer linear kernels are more adapted in the ETT dataset than our Transformer and LSTM kernels.
-We remark that K-U-Net Linear_1000 is outstanding in the Electricity (S) dataset and thus selected for full examination.
-A.5 Result of selected variants
-In Table 5 6 7, we note selected models that are examined prudently with 5 runs in a fine-tuned context and measures with average MSE and MAE.
-We include the Table 9 showing that K-U-Net has a small standard deviation error across 5 random runs.
-A.6 Result of univariate time series forecasting
-The experiment of univariate time series forecasting task predicts the target feature is the ”oil temperature” on ETT datasets.
-We remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3).
-It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ in the Univariate setting compared to PatchTST.
-Appendix B Ablation Study
-For studying the roles of skip connection in K-U-Net, Random Erase, and Chanel Independence, we use the Minimum Top5 average MSE (Min Top5MMSE) to measure the extreme performance and the Mean Top5MMSE to measure the average performance.
-For instance Normalisation, we only show the comparison with Mean Normalization with the Traffic dataset because the improvement is evident.
-Table 5: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm.
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold.
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_0000
-(720) K-U-Net
-Linear_1100
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.372 0.393 0.377 0.401 0.370 0.407 0.370 0.393 0.370 0.400
-192 0.405 0.414 0.413 0.424 0.404 0.433 0.404 0.414 0.413 0.429
-336 0.431 0.430 0.444 0.444 0.420 0.445 0.420 0.430 0.422 0.440
-720 0.438 0.454 0.457 0.471 0.482 0.490 0.438 0.454 0.447 0.468
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_0000
-(720) K-U-Net
-Linear_0001
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh2
-96 0.273 0.335 0.271 0.336 0.286 0.342 0.271 0.335 0.274 0.337
-192 0.336 0.377 0.332 0.377 0.354 0.384 0.332 0.377 0.339 0.379
-336 0.360 0.401 0.357 0.403 0.388 0.413 0.357 0.401 0.329 0.384
-720 0.395 0.435 0.394 0.442 0.420 0.447 0.394 0.435 0.379 0.422
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_0000
-(720) K-U-Net
-Linear_1000
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1
-96 0.301 0.343 0.310 0.351 0.286 0.342 0.286 0.342 0.290 0.342
-192 0.336 0.363 0.338 0.368 0.330 0.371 0.330 0.363 0.332 0.369
-336 0.371 0.384 0.367 0.385 0.360 0.393 0.360 0.384 0.366 0.392
-720 0.427 0.416 0.417 0.412 0.405 0.420 0.405 0.412 0.416 0.420
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_0000
-(720) K-U-Net
-Linear_0110
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm2
-96 0.164 0.253 0.161 0.252 0.165 0.256 0.161 0.252 0.165 0.255
-192 0.219 0.290 0.216 0.290 0.225 0.296 0.216 0.290 0.220 0.292
-336 0.273 0.326 0.268 0.326 0.275 0.332 0.268 0.326 0.274 0.329
-720 0.367 0.383 0.350 0.379 0.343 0.380 0.343 0.379 0.362 0.385
-Table 6: Fine-tuned multivariate time series forecasting results with Kernel U-Net on datasets Electricity, Weather and Traffic.
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold.
-Method K-U-Net
-Linear_0000
-(720) K-U-Net
-Linear_1110
-(720) K-U-Net
-Linear_1000
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Electricity
-96 0.138 0.237 0.134 0.232 0.129 0.226 0.129 0.226 0.129 0.222
-192 0.153 0.250 0.152 0.249 0.147 0.244 0.147 0.244 0.147 0.240
-336 0.171 0.269 0.167 0.264 0.163 0.261 0.163 0.261 0.163 0.259
-720 0.213 0.305 0.203 0.300 0.197 0.292 0.197 0.292 0.197 0.290
-Method K-U-Net
-Linear_0000
-(720) K-U-Net
-Transf_0001
-(720) K-U-Net
-LSTM_0001
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Traffic
-96 0.412 0.280 0.399 0.262 0.358 0.253 0.358 0.253 0.360 0.249
-192 0.425 0.287 0.415 0.278 0.373 0.262 0.373 0.262 0.379 0.250
-336 0.438 0.294 0.424 0.295 0.390 0.271 0.390 0.271 0.392 0.264
-720 0.458 0.314 0.465 0.309 0.430 0.292 0.430 0.292 0.432 0.286
-Method K-U-Net
-Linear_0110
-(720) K-U-Net
-Transf_0001
-(720) K-U-Net
-LSTM_0100
-(720) MIN PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-Weather
-96 0.146 0.201 0.146 0.200 0.142 0.195 0.142 0.195 0.149 0.198
-192 0.190 0.245 0.192 0.246 0.188 0.244 0.188 0.244 0.194 0.241
-336 0.241 0.284 0.242 0.287 0.242 0.286 0.241 0.284 0.245 0.282
-720 0.310 0.333 0.310 0.336 0.316 0.341 0.310 0.333 0.314 0.334
-B.1 Instance Normalisation and Mean Normalisation
-Instance normalization contains a step of substruction with the average value of the input sequence and a division of its standard deviation.
-It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model.
-Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises into the dataset.
-We employed the mean normalization as the preprocessing step for most experiments except those with instance normalization on traffic datasets.
-The comparison in Table 8 shows that Instance Normalization improves the MSE by 3.89% for a K-U-Net LSTM_0001 on traffic datasets.
-Table 7: Fine-tuned univariate time series forecasting results with Kernel U-Net on datasets ETTh and ETTm.
-The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets.
-We note the best results in bold.
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_1111
-(336) MIN PatchTST Nlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-96 0.053 0.177 0.054 0.182 0.053 0.177 0.055 0.179 0.053 0.177
-192 0.069 0.205 0.079 0.219 0.069 0.205 0.071 0.215 0.069 0.204
-336 0.076 0.222 0.076 0.221 0.076 0.221 0.076 0.22 0.081 0.226
-720 0.077 0.224 0.078 0.226 0.077 0.224 0.087 0.232 0.08 0.226
-Method K-U-Net
-Linear_0110
-(336) K-U-Net
-Linear_1100
-(336) MIN PatchTST Nlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh2
-96 0.114 0.267 0.121 0.272 0.114 0.267 0.129 0.282 0.129 0.278
-192 0.146 0.306 0.161 0.317 0.146 0.306 0.168 0.328 0.169 0.324
-336 0.157 0.321 0.178 0.340 0.157 0.321 0.171 0.336 0.194 0.355
-720 0.192 0.355 0.182 0.347 0.182 0.347 0.223 0.38 0.225 0.381
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_0000
-(720) MIN PatchTST Nlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1
-96 0.026 0.121 0.026 0.122 0.026 0.121 0.026 0.121 0.026 0.122
-192 0.038 0.149 0.039 0.150 0.038 0.149 0.039 0.15 0.039 0.149
-336 0.051 0.173 0.051 0.174 0.051 0.173 0.053 0.173 0.052 0.172
-720 0.072 0.206 0.068 0.203 0.068 0.203 0.073 0.206 0.073 0.207
-Method K-U-Net
-Linear_0000
-(336) K-U-Net
-Linear_1000
-(720) MIN PatchTST Nlinear
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm2
-96 0.063 0.182 0.062 0.183 0.062 0.182 0.065 0.187 0.063 0.182
-192 0.090 0.224 0.089 0.227 0.089 0.224 0.093 0.231 0.09 0.223
-336 0.116 0.259 0.114 0.263 0.114 0.259 0.121 0.266 0.117 0.259
-720 0.170 0.319 0.156 0.312 0.156 0.312 0.171 0.322 0.17 0.318
-Table 8: Comparison of Mean Normalization(+MN) and Instance Normalization (+IN) on Traffic dataset. We note the best results in bold..
-Methods K-U-Net (+MN)
-LSTM_0001
-(720) K-U-Net (+IN)
-LSTM_0001
-Metric MSE MAE MSE MAE
-Traffic
-96 0.375 0.264 0.360 0.252
-192 0.393 0.275 0.377 0.263
-336 0.406 0.280 0.389 0.270
-720 0.442 0.296 0.427 0.290
-Table 9: MSE and standard deviation of K-U-Net across 5 random runs.
-Method K-U-Net
-LSTM_0001
-(720) Method K-U-Net
-LSTM_0100
-(720) Method K-U-Net
-Linear_1000
-Metric MSE Metric MSE Metric MSE
-Traffic
-96 0.35761 ±0.00042
-Weather
-96 0.14200 ±0.00016
-ETTm1
-96 0.28600 ±0.00056
-192 0.37266 ±0.00056 192 0.18782 ±0.00078 192 0.32996 ±0.00204
-336 0.39030 ±0.00135 336 0.24164 ±0.00139 336 0.35979 ±0.00092
-720 0.42970 ±0.00078 720 0.31646 ±0.00198 720 0.40539 ±0.00128
-B.2 Skip Connection,
-Skip Connection in the U-Net is used for passing information from the encoder to the same level on its symmetric decoder.
-As the size of the hidden vector is reduced during passage in the encoder, skip connection helps keep the low-level features and allows the decoder to mix with high-level information.
-We conduct experiments to understand the role of skip connection in the K-U-Net.
-The result in Table 4 shows that removing the skip connection degrades the Min Top5MMSE score for $0.77\%$ on average and especially $3.19\%$ on the ETTh1 dataset.
-Additionally, it degrades $2.30\%$ Mean Top5MMSE on average and especially $5.42\%$ on the ETTh1 dataset.
-B.3 Comparison of Chanel Independent and Chanel Mixing
-Chanel Independent was an inherited choice from Zeng et al.
-[2023].
-It avoids introducing noises that degrade the performance across different features.
-In this experiment, we compare the channel-independent(CI) strategy and the channel-mixing strategy.
-In K-U-Net, setting input and output dimension to 1 implies a channel-independent strategy, while setting to feather size (7 for the ETT dataset) means a channel-mixing strategy.
-As We observe in Table 4, removing CI degrades the Min Top MMSE score for $21.45\%$ and the Min Top5MMSE score for $23.72\%$.
-B.4 Comparison of Random Erasing
-Random Erasing augments the dataset by masking partially the input.
-In this experiment, we studied the role of random erasing in time series forecasting settings.
-We observe in Table 4 that removing Random Erasing degrades the Min Top5MMSE score for $0.2\%~{}0.32\%$, which is not as evident as the Skip Connection and Channel Independent strategy.
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Weather
-(f) Weather(S)
-(g) Electricity(S)
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Weather
-(f) Weather(S)
-(g) Electricity(S)
-(h) Traffic(S)
-Figure 9: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on validation and test set.
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Weather
-(f) Weather(S)
-(g) Electricity(S)
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Weather
-(f) Weather(S)
-(g) Electricity(S)
-(h) Traffic(S)
-Figure 10: Search result of K-U-Net with Multi-layer Linear kernels on validation and test set.